### v3

# A computable version of Hall’s Harem Theorem and Geometric von Neumann Conjecture

Abstract.
We prove a version of Hall’s Harem Theorem where the matching realizes a function with some special properties of its cycles. Using this approach we also prove a new computable version of Hall’s Harem theorem. We apply it to non-amenable computable coarse spaces.
1. Introduction
Analysis of classical mathematical topics from the point of view of complexity of various types is one of the major trends of modern mathematical logic. Amenability is essentially fruitful from this point of view (see [Kec91], [BK96], [KPT05], [MU16], [CK18], [HPP08], [HKP22] and [HKP]). Our research belongs to computable amenability. This is a topic where computable versions of fundamentals of amenability are studied, see papers of M. Cavaleri [Cav17, Cav18], N. Moryakov [Mor18] and the author (together with A. Ivanov) [DI22b, DI22a]. In the present paper we are particularly interested in von Neumann conjecture: the statement that each non-amenable group contains a subgroup isomorphic to $\mathbb{F}_{2}$.
Let $X$ be a set and let $G$ be a group which acts on $X$ by permutations. Then the $G$-space $(G,X)$ is called amenable if there exists a mean $m:\ell^{\infty}(X)\rightarrow\mathbb{R}$ which is invariant under the action of $G$. In particular, if $X=G$, then the group $G$ is called amenable. This definition is due to to J. von Neumann [Neu29]. In 1938, Tarski proved his famous Tarski Alternative Theorem [Tar38], stating that a $G$-space $X$ is amenable if and only if $X$ does not admit a $G$-paradoxical decomposition. Some computable versions of the Tarski Alternative Theorem have been found by the author together with A. Ivanov in [DI22a]. Roughly speaking effective paradoxical deocmpositions of $G$-spaces of computable transformations were introduced and it was proved that their existence is equivalent to non-amenability.
In [Neu29] J. von Neumann proved that amenable groups cannot contain a subgroup isomorphic to the two generated free group $\mathbb{F}_{2}$. The question whether the converse holds i.e. whether each non-amenable group contains a subgroup isomorphic to $\mathbb{F}_{2}$, became popularized as the von Neumann Conjecture in the late 50s (see e.g. [Day57]). It was answered negatively with the construction of Tarski monster group by Olshanskii in 1980 [Ol’81]. However, in 1999 Whyte [Why99] proved so called the geometric version of this conjecture showing existence of $d$-regular forests in non-amenable metric spaces. The latter result is now called Geometric von Neumann Conjecture. The name corresponds to an easy corollary that so called the wobbling group of the space has a subgroup isomorphic to $\mathbb{F}_{2}$ with a semi-regular action. In 2018 Schneider [Sch18] generalized Whyte’s theorem and the corollary to coarse spaces.
The main results of our paper are:
  • computable versions of Schneider’s theorem and the corollary about wobbling groups,
  • a computable version of the Hall Harem Theorem.
The Hall Harem Theorem is a generalization of Hall’s marriage theorem, see [CSC10, Theorem H.4.2.]. We consider it together with geometric von Neumann conjecture just because of the well-known fact that aforementioned classical results can be obtained by applications of this theorem.
1.1. A computable version of Scheider’s theorem
The theorem of Whyte mentioned above is as follows.
Theorem 1.1 (Geometric Von Neumann Conjecture, [Why99]).
A uniformly discrete metric space of uniformly bounded geometry is non-amenable if and only if it admits a partition whose pieces are uniformly Lipschitz embbeded copies of the $4$-regular tree.
Since we concentrate on Schneider’s generalization of this theorem [Sch18] we skip thorough explanation of the terminology used by Whyte. Schneider’s theorem [Sch18] concerns non-amenable coarse spaces. We will introduce them to the reader. This would give rough understanding of the topic of Geometric von Neumann Conjecture.
Let us recall some terminology from coarse geometry (according to [Sch18]). For a relation $E\subseteq X\times X$ on a set $X$ and $x\in X,A\subseteq X$, let
$E[x]:=\{y\in x|(x,y)\in E\},$
and
$E[A]:=\bigcup\{E[x]|x\in A\}.$
Furthermore, we denote by $\Gamma(E)$ the graph associated with the relation $E$, i.e. $\Gamma(E)=(X,E)$.
Definition 1.2.
A coarse space is a pair $(X,\mathcal{E})$ consisting of a set $X$, and a collection $\mathcal{E}$ of subsets of $X\times X$ (called entourages) such that:
  • the diagonal $\Delta_{X}\in\mathcal{E}$;
  • if $F\subseteq E\in\mathcal{E}$, then also $F\in\mathcal{E}$;
  • if $E,F\in\mathcal{E}$ then $E\cup F,E^{-1},E\circ F\in\mathcal{E}$.
Definition 1.3.
A coarse space $(X,\mathcal{E})$ is said to have bounded geometry if for each $E\in\mathcal{E}$ and every $x\in X$ the set $E[x]$ is finite.
Note that metric spaces are examples of coarse spaces: for a given metric space $X$, a coarse space $(X,\mathcal{E})$ is obtained by setting:
:=$\mathcal{E}:=\{E\subseteq X\times X|\sup\{d(x,y)|(x,y)\in E\}\leq\infty\}.$
Definition 1.4.
A coarse space $(X,\mathcal{E})$ of bounded geometry is called amenable if for every $\theta>1$ and every $E\in\mathcal{E}$ there exists a non-empty finite $F\subseteq X$ such that $|E[F]|\leq\theta|F|.$
Theorem 1.5.
([Sch18, Theorem 2.2]) Let $d\geq 3$. A coarse space $(X,\mathcal{E})$ of bounded geometry is not amenable if and only if there is $E\in\mathcal{E}$ such that $\Gamma(E)$ is a $d$-regular forest.
To see why this theorem generalizes the theorem of Whyte, note that connected components of $\Gamma(E)$ can be viewed as pieces of the partition from Whyte’s theorem.
In the proof of his result, Schneider takes a symmetric entourage $E$ such that $|E(F)|\geq(d-1)|F|$ for all finite $F\subset X$. Such an entourage exists by the fact that when $(X,\mathcal{E})$ is not amenable, then for every finite $d$ there is a symmetric entourage such that $|E(F)|\geq d|F|$ for each finite $F\subseteq X$ (see discussion before Proposition 2.1 in [Sch18]). Now the bipartite graph $\Gamma:=(X,X,R)$ where $R:=E\setminus\Delta_{X}$, satisfies Hall’s Harem condition for $d-1$:
for all finite subsets $Y\subset X$ the following inequality holds:
|$|\{x\,|\,\exists y\in Y(x,y)\in R\}|\geq(d-1)|Y|$.
It follows by Hall’s harem theorem that there exists an entourage which graph is $d$-regular. Moreover, properties of entourages allows us to make it cycle-free without losing $d$-regularity.
In our paper we effectivize this scheme. The final goal is formulated as follows. (In the formulation the term a highly computable graph just means a computable locally finite graph where the valency of each vertex can be effectively recognized.)
Theorem A.
Let $d\geq 3$. Let a coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that for any finite $F\subseteq\mathbb{N}$ we have $|E[F]|\geq(d+2)|F|$. Then there exists a computable $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a $d$-regular forest. Moreover, there exist an algorithm which for any $m,n\in\mathbb{N}$ recognizes if $m$ and $n$ are in the same connected component of $\Gamma(E^{\prime})$.
To see that this statement is a computable version of Schneider’s result we remind the reader that as we already mentioned when $(X,\mathcal{E})$ is not amenable, then for every finite $d$ there is a symmetric entourage such that $|E(F)|\geq d|F|$ for every finite $F\subseteq X$.
We now translate Theorem A into a statement about wobbling groups. Let us recall that given a coarse space $(X,\mathcal{E})$ its wobbling group is
$\mathcal{W}(X,\mathcal{E}):=\{\alpha\in Sym(X)|\{(x,\alpha(x))|x\in X\}\in\mathcal{E}\}.$
Wobbling groups have attracted growing attention in recent years [JM13], [JdlS15], [Sch18]. One of motivating points is a possibility of a nice reformulation of Whyte’s result in terms of semi-regular subgroups. Recall that a subgroup $G\leq Sym(X)$ is said to be semi-regular if no non-identity element of $G$ has a fixed point in $X$. Theorem 6.1 in [Why99] states that a uniformly discrete metric space $X$ of uniformly bounded geometry is non-amenable if and only if $\mathsf{F}_{2}$ is isomorphic to a semi-regular subgroup of $\mathcal{W}(X)$. Corollary 2.3 of [Sch18] states the same for coarse spaces. Note that these formulations look like von Neumann conjecture!
The following theorem is a computable version of Corollary 2.3 of [Sch18].
Theorem B.
Let the coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that $|E[F]|\geq 6|F|$ for all finite $F\subset\mathbb{N}$. Then there are two computable permutations $\sigma,\pi\in Sym(\mathbb{N})$ such that $\langle\sigma,\pi\rangle$ is a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$.
1.2. Hall’s harem theorem
The Schneider’s scheme presented above is based on an application of the Hall Harem theorem. The latter describes a condition which is equivalent to existence of a perfect $(1,k)$-matching of a bipartite graph, for example see Section III.3 in [Bol79], Theorem 2.4.2 in [LP86] or Theorem H.4.2 in [CSC10].
In order to obtain a computable version of Schneider’s theorem some special form of Hall’s harem theorem is required. Assume that the graph is of the form $(\mathbb{N},\mathbb{N},E)$, where $E$ denotes the set of edges between natural numbers. Any matching clearly realizes a $k$ to $1$ function, say $f:\mathbb{N}\rightarrow\mathbb{N}$. The main concern of our paper are properties which can be added to such a function. In particular we are especially interested in fast reaching cycles of $f$. This will be formalized below as the property of controlled sizes of cycles. As a result the matching obtained in this way, is realized by a computable subset of $\mathbb{N}^{2}$ with the additional property that there is an algorithm which decides whether two natural numbers $m$ and $n$ belong to the same connected component of $f$.
The special form of Hall’s harem theorem mentioned in the previous paragraph is the heart of our paper and the proof of it (as we evaluate) is 75 per cent of the text. In Sections 2 and 3 (a short one) we give an elementary introduction to both graph theoretic and computability theoretic issues. In Section 4 we formulate the computable version of Hall’s harem theorem, which is our goal. There we also describe the strategy of the proof. It is worth noting here that the proof is in the standard style of building of computably enumerable sets as it is usually done in computability theory. We worn the reader that the construction is technical and difficult. A kind of priority arguments is used in it. The proof of it is given in Sections 3 - 6. In Section 7 we connect this material with computable amenability.
The preprint [Dud21] contains the material of the present paper. It also contains the simplified version of Hall’s harem theorem as above where all computability elements are ruled out. We should mention that these two versions are incomparable, but complement each other so that naturally form two parts of a research project.
Remark 1.6.
In the final remark of this introduction we mention other investigations concerning Hall’s theorems in the context of computability theory. The topic was initiated in papers of Manaster and Rosenstain [Man72] and Kierstead [Kie83]. The latter paper was our starting point. Hall’s theorems have become quite popular in complexity theory. For example see [PZ80], [IRT78].
From the point of view of reverse mathematics Hall’s theorems are studied in [FJLM22], [Hir90] and [HH22].
2. Preliminaries
2.1. General preliminaries
To introduce the reader to the subject we recall Hall’s Harem theorem. We begin with some necessary definitions from graph theory. We mostly follow the notation of [CSC10]. A graph $\Gamma=(\mathbf{\Gamma},E)$ is called a bipartite graph if the set of vertices $\mathbf{\Gamma}$ is partitioned into sets $U$ and $V$ in such a way, that the set of edges $E$ is a subset of $U\times V$. We denote such a bipartite graph by $\Gamma=(U,V,E)$.
From now on we concentrate on bipartite graphs. Note that although the definitions below concern this case, they usually have obvious extensions to all ordinary graphs.
A subgraph of $\Gamma$ is a triple $\Gamma^{\prime}=(U^{\prime},V^{\prime},E^{\prime})$ with $U^{\prime}\subseteq U$, $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$. When $\Gamma^{\prime}$ is such a subgraph but only the sets of its vertices are specified (i.e. for example $\Gamma^{\prime}=(U^{\prime},V^{\prime})$), this means that $E^{\prime}=E\cap(U^{\prime}\times V^{\prime})$. Then we say that the subgraph $\Gamma^{\prime}$ is induced in $\Gamma$ by the set of vertices.
Let $\Gamma=(U,V,E)$. We will say that an edge $(u,v)$ is incident to vertices $u$ and $v$. In this case we say that $u$ and $v$ are adjacent. When two edges $(u,v),(u^{\prime},v^{\prime})\in E$ have a common incident vertex we say that $(u,v),(u^{\prime},v^{\prime})$ are also adjacent. A sequence $x_{1},x_{2},\ldots,x_{n}$ is called a path if each pair $x_{i},x_{i+1}$ is adjacent, $1\leq i<n$.
Below we will denote the set of vertices $\mathbf{\Gamma}$ by the same letter with the graph as a structure, i.e. $\Gamma$. Given a vertex $x\in\Gamma$ the neighbourhood of $x$ is a set
$N_{\Gamma}(x)=\{y\in\Gamma:(x,y)\in E\}.$
For subsets $X\subset U$ and $Y\subset V$, we define the neighbourhood $N_{\Gamma}(X)$ of $X$ and the neighbourhood $N_{\Gamma}(Y)$ of $Y$ by
$N_{\Gamma}(X)=\bigcup\limits_{x\in X}N_{\Gamma}(x)\,\,\text{ and }\,\,N_{\Gamma}(Y)=\bigcup\limits_{y\in Y}N_{\Gamma}(y).$
We drop the subscript $\Gamma$ if it is clear from the context.
For a given vertex $v$ a star of $v$ is a subgraph $S=(\{v\}\cup N_{\Gamma}(v),E^{\prime})$ of $\Gamma$, with $E^{\prime}=((\{v\}\cup N_{\Gamma}(v))\times(\{v\}\cup N_{\Gamma}(v)))\cap E$. A $(1,k)$-fan is a subset of $E$ consisting of $k$ edges incident to some vertex $u\in U$. We say that $u$ is the root of the fan, and when $(u,v)$ belongs to the fan we call $v$ a leaf of it.
Definition 2.1.
An $(1,k)$-matching from $U$ to $V$ is a collection $M$ of pairwise disjoint $(1,k)$-fans.
The $(1,k)$-matching $M$ is called left perfect (resp. right perfect) if each vertex from $U$ is a root of a fan from $M$ (resp. each vertex from $V$ belongs to exactly one fan of $M$). The $(1,k)$-matching $M$ is called perfect if it is both left and right perfect.
We often view an $(1,k)$-matching as a bipartite graph $M$ where the fan of $u\in U$ is the $M$-star of $u$, i.e. a graph consisting of the set of all vertices adjacent to $u$ in $M$ and all edges incident to $u$ in $M$. We emphasize that a perfect $(1,k)$-matching from $U$ to $V$ is a set $M\subset E$ satisfying following conditions:
  1. (1)
    for each vertex $u\in U$ there exists exactly $k$ vertices $v_{1},\ldots,v_{k}\in V$ such that $(u,v_{1}),\ldots,(u,v_{k})\in M$;
  2. (2)
    for all $v\in V$ there is an unique vertex $u\in U$ such that $(u,v)\in M$.
Originally the Hall’s Marriage Theorem (see e.g. [Die97, Theorem 2.1.2], [Bol79, Section III.3], [LP86, Theorem 2.4.2.]) provides us a condition for existence of left perfect $(1,1)$-matching in a finite bipartite graph. We are interested in so called Hall’s Harem theorem, which is a generalization of Hall’s Marriage theorem to the case of perfect $(1,k)$-matchings for the locally finite infinite graphs.
One says that graph $\Gamma=(V,E)$ is locally finite if for all vertices $x\in\Gamma$, the neighbourhood $N(x)$ is finite. Note that if $\Gamma$ is locally finite, then $N(X)$ is finite for any finite $X\subset V$.
Theorem 2.2 (Hall’s Harem theorem).
[CSC10, Theorem H.4.2.] Let $\Gamma=(U,V,E)$ be a locally finite graph and let $k\in\mathbb{N},\;k\geq 1$. The following conditions are equivalent:
  1. (i)
    For all finite subsets $X\subset U$, $Y\subset V$ the following inequalities holds: $|N(X)|\geq k|X|$, $|N(Y)|\geq\frac{1}{k}|Y|$.
  2. (ii)
    $\Gamma$ has a perfect $(1,k)$-matching.
The first condition in this formulation is known as Hall’s $k$-harem condition.
It is a crucial fact that the theorem holds for locally finite infinite graphs. At this point we inform the reader that the statement of the Hall Marriage Theorem does not work for infinite graphs in general (see e.g. [Die97, S.2 Ex.6]). Nevertheless, there are versions of this theorem for graphs of any cardinality [Aha88].
2.2. Reflections
Throughout the paper, $d$ is a natural number greater than $1$. When $\Gamma=(U,V,E)$ is a bipartite graph, we always assume that $V\subseteq U\subseteq\mathbb{N}$, i.e. $V$ is a subset of the right copy of $U$.
The following notation substantially simplifies the presentation. For any vertex $v\in V$ there exist a vertex from $U$ which is a copy of $v$ (i.e. the same natural number), we denote it by $u_{v}$. If a vertex $u\in U$ has the copy in $V$ then we denote this copy by $v_{u}$.
Definition 2.3.
The graph $\Gamma=(U,V,E)$ is called $U$-reflected if $V$ is a subset of the right copy of $U$ and for every edge $(u,v)\in E$ with $v_{u}\in V$ the edge $(u_{v},v_{u})$ is in $E$ too. If additionally $V$ is a right copy of $U$ then $\Gamma$ is called a fully reflected bipartite graph.
The Theorem 2.5 below states that in the case of fully reflected bipartite graphs the Hall’s harem condition allows us to force some additional properties at the expense of obtaining a perfect $(1,(d-1))$-matching instead of a $(1,d)$-matching. This is an intermediate statement that avoids computability issues as much as it is possible, the proof of this theorem is available in [Dud21]. We will now give necessary details.
2.3. Controlled sizes of cycles.
Let $f$ be a function. If for some $i\neq 0$ we have $f^{i}(u)=u$ then we will say that $u$ is a periodic point. For such $u$ and the smallest $i\neq 0$ with $f^{i}(u)=u$, we say that $\{u,f(u),\ldots,f^{i-1}(u)\}$ is a cycle of $f$. Any $(1,(d-1))$ matching can be considered as a $(d-1)$ to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$. Moreover, if this matching is perfect, such a function $f$ is total and surjective. We roughly want to show that given a fully reflected bipartite graph satisfying Hall’s $d$-harem condition, there is a perfect $(1,(d-1))$ matching $f:\mathbb{N}\rightarrow\mathbb{N}$, such that for each $u$ there exist $i\geq 0$ such that $f^{i}(u)$ is a periodic point.
Definition 2.4.
Let $f:\mathbb{N}\rightarrow\mathbb{N}$ be a $(d-1)$ to $1$ function. We say that $f$ has controlled sizes of its cycles if each of the following conditions holds:
  1. (i)
    $f^{2}(1)=1$;
  2. (ii)
    if $n\geq 2$ and $f^{i}(n)=n$ then $i\leq n$;
  3. (iii)
    if $n\geq 2$ and for all $i\leq n$ we have $f^{i}(n)\neq n$ then there exist $k\leq 2n$ and $l\leq n$ such that $f^{k+l}(n)=f^{k}(n)$.
The following theorem is proved in [Dud21].
Theorem 2.5.
Let $\Gamma=(U,V,E)$ be a locally finite bipartite graph such that:
  • both $U$ and $V$ are identified with $\mathbb{N}\setminus\{0\}$,
  • $E$ does not contain edges of the form $(u,v_{u})$,
  • $\Gamma$ is fully reflected,
  • $\Gamma$ satisfies Hall’s $d$-harem condition.
Then there exist a perfect $(1,d-1)$-matching of $\Gamma$, which realizes a $(d-1)$ to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$ with controlled sizes of its cycles.
3. Computable versions of Hall’s $d$-harem theorem.
Before we state a computable version of Theorem 2.5 we describe our approach to computability of classical Hall’s $d$-harem theorem from [DI22a].
Let us consider the following definition. It is standard in computability theory.
Definition 3.1.
A graph $\Gamma$ with the set of vertices $\Gamma$ is computable if there exists a bijective function $\nu:\mathbb{N}\rightarrow\Gamma$ such that the set
$R:=\{(i,j):(\nu(i),\nu(j))\in E\}$
is computable.
A bipartite graph $\Gamma=(U,V,E)$ (with $\Gamma=U\,\dot{\cup}\,V$) is computably bipartite if $\Gamma$ is computable and the set of $\nu$-numbers of $U$ is computable.
To simplify the matter, in the case of bipartite graphs, from now on we will always identify $U$ and $V$ with computable subsets of $\mathbb{N}$ so that $\nu$ is the identity map. Further, admitting that $U\cap V\not=\emptyset$ we distinguish these sets saying that $U$ is taken from the left copy of $\mathbb{N}$ but $V$ is taken from the right one. In particular we often consider graphs where $U=V=\mathbb{N}$.
To formulate a computable version of Hall’s Harem theorem we need the following definition of Kierstead [Kie83].
Definition 3.2.
A locally finite computable graph $\Gamma$ is called highly computable if the function $g(n)=|N_{\Gamma}(\nu(n))|$, $n\in\mathbb{N}$, is a computable function $g:\mathbb{N}\rightarrow\mathbb{N}$.
It is worth noting here that this definition easily implies the existence of an algorithm which for every $m$ and $n$ computes the $n$-ball of $\nu(m)$ in $\Gamma$.
Manaster and Rosenstein [Man72] showed that there is a computably bipartite, highly computable graph that satisfies Hall’s condition for a bipartite graph to have a $(1,1)$-matching, but has no computable $(1,1)$-matching.
Kierstead [Kie83] found the natural extension of Hall’s condition, called computable expanding Hall’s condition, which allows us to construct computable $(1,1)$-matchings.
Following [DI22a], in this paper we focus on computable perfect $(1,k)$-matchings.
Definition 3.3.
Let $\Gamma=(U,V,E)$ be a computably bipartite graph. A perfect $(1,k)$-matching $M$ from $U$ to $V$ is called a computable perfect $(1,k)$-matching if $M$ is a computable set of pairs.
Observe that computable perfectness exactly means that there is an algorithm which
  • for each $i\in U$, finds the tuple $(i_{1},i_{2},\ldots,i_{k})$ such that $(i,i_{j})\in M$, for all $j=1,2,\ldots,k$;
  • when $i\in V$ it finds $i^{\prime}\in U$ such that $(i^{\prime},i)\in M$.
In [DI22a] we introduced the following condition, which implies the existence of computable perfect $(1,k)$-matchings in highly computable bipartite graphs. For $k=1$ it was formulated earlier in [Kie83].
Definition 3.4.
A bipartite graph $\Gamma=(U,V,E)$ satisfies the computable expanding Hall’s harem condition with respect to $k$ (denoted $c.e.H.h.c.(k)$), if and only if there is a total, computable function $h:\mathbb{N}\rightarrow\mathbb{N}$ such that:
  • $h(0)=0$
  • for all finite sets $X\subset U$, the inequality $h(n)\leq|X|$ implies $n\leq|N(X)|-k|X|$
  • for all finite sets $Y\subset V$, the inequality $h(n)\leq|Y|$ implies $n\leq|N(Y)|-\frac{1}{k}|Y|$.
In [DI22a] we have proven the following.
Theorem 3.5.
If $\Gamma=(U,V,E)$ is a highly computable bipartite graph satisfying the $c.e.H.h.c.(k)$, then $\Gamma$ has a computable perfect $(1,k)$-matching.
4. A computable version of Hall’s harem Theorem with cycles, preliminaries
We assume that $\Gamma=(U,V,E)$ is a bipartite graph, such that $V\subseteq U\subseteq\mathbb{N}$, i.e. $V$ is a subset of the right copy of $U$.
4.1. Main Theorem and the aim of the construction
In the proof of Theorem A we need a computable perfect $(1,(d-1))$ matching which realizes a function $f:\mathbb{N}\rightarrow\mathbb{N}$, such that the following conditions are satisfied:
  1. (I)
    $\{n\,|\,n\text{ is a periodic point}\}$ is computable;
  2. (II)
    $\{(n,m)\,|\,n\text{ is a periodic point },m\text{ is in the cycle of }n\}$ is computable;
  3. (III)
    the set $\{(n,m)\,|\,f^{m}(n)\text{ is a periodic point}\}$ is computable and its first coordinates cover $\mathbb{N}\setminus\{0\}$.
Moreover, the matching has to be constructed by a uniform algorithm.
In fact it is easy to see that a computable $(d-1)$ to $1$ function with controlled sizes of its cycles (Definition 2.4) satisfies conditions (I)–(III).
The following theorem is a kind of amalgamation of Theorem 2.5 and Theorem 3.5.
Theorem 4.1 (Main Theorem).
Let $\Gamma=(U,V,E)$ be a highly computable bipartite graph, such that:
  • both $U$ and $V$ are identified with $\mathbb{N}\setminus\{0\}$
  • $E$ does not contain edges of the form $(u,v_{u})$
  • $\Gamma$ is fully reflected,
  • $\Gamma$ satisfies the $c.e.H.h.c.(d)$.
Then there exist a computable perfect $(1,d-1)$-matching of $\Gamma$, which realizes a computable $(d-1)$ to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$ with controlled sizes of its cycles.
Main Theorem is proved in the section 8. The next sections focus on a construction used in this proof.
4.2. Notation used in the construction
Since this construction is highly technical, we start with a list of the notation. We do not insist on a thorough inspection of it. In the beginning a hasty view will suffice.
  • $M$ is a perfect matching that we construct.
  • $M_{n-1}$ is a set of $(1,d-1)$-fans added to $M$ at the end of the $n$-th step. Thus $M=\bigcup\limits_{n=1}^{\infty}M_{n-1}$.
  • $\Gamma^{(0)}=(U^{(0)},V^{(0)},E^{(0)})$ is the original graph $\Gamma$.
  • $U^{(n)}:=U^{(n-1)}\setminus\{u\in U^{(n-1)}:\exists v\in V^{(n-1)},(u,v)\in M_{n-1}\}$.
  • $V^{(n)}:=V^{(n-1)}\setminus\{v\in V^{(n-1)}:\exists u\in U^{(n-1)},(u,v)\in M_{n-1}\}$.
  • $\Gamma^{(n)}=(U^{(n)},V^{(n)})$. We will see that $\Gamma^{(n)}$ is $U^{(n)}$-reflected.
  • After the $n$-th step we obtain decompositions $U^{(n)}=U^{(n)\star}\,\dot{\cup}\,U^{(n)\perp}$ and $V^{(n)}=V^{(n)\star}\,\dot{\cup}\,V^{(n)\perp}$, where we say that $U^{(n)\perp}$ consists of elements from $U^{(n)}$ which might spoil Hall’s $d$-harem condition for $\Gamma^{(n)}$.
  • Put $U^{(0)\perp}=\emptyset$ and $V^{(0)\perp}=\emptyset$ (since $\Gamma^{(0)}$ is $\Gamma$, i.e. it satisfies Hall’s $d$-harem condition).
  • $\Gamma^{(n)\perp}$ is a graph with the sets of vertices $(U^{(n)\perp},V^{(n)\perp})$ and the set of edges corresponding to $(1,d-1)$-fans with roots denoted by $u^{\perp}_{i}\in U^{(n)\perp}$.
  • When $U^{(n)\perp}\setminus U^{(n-1)\perp}$ is not empty, $U^{(n)\perp}\setminus U^{(n-1)\perp}=\{u^{\perp}_{n-1}\}$ and $V^{(n)\perp}\setminus V^{(n-1)\perp}$ consists of leaves $\{v^{\perp}_{n-1,i}:1\leq i\leq d-1\}$.
  • $\Gamma^{(n)\star}:=(U^{(n)}\setminus U^{(n)\perp},V^{(n)}\setminus V^{(n)\perp})$. We will see that $\Gamma^{(n)\star}$ satisfies Hall’s harem condition.
  • During the construction of $M_{n}$ we will define fans $M^{j}_{n}$, $j\leq n+1$. The graph $M_{n}$ is the union of them.
  • $M^{j}_{n}$ is a fan consisting of edges denoted by $\{(u^{j}_{n},v^{j}_{n,i}):i\leq d-1\}$.
  • $u_{n}$ is a starting vertex of the $n$-th step, it is also denoted by $u^{0}_{n}$.
  • For any subgraph $\Gamma^{\prime}$ of $\Gamma^{(n)}$ by $\Gamma^{\prime}(-u^{0}_{n},\ldots,-u^{j}_{n})$ we denote the graph obtained from $\Gamma^{\prime}$ by removal of the $(1,d-1)$-fans of $M_{n}$ with roots $u^{0}_{n},\ldots,u^{j}_{n}$.
  • For any subgraph $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ of $\Gamma^{(n)}$ and any $u_{j}^{\perp}\in U^{(n)\perp}$ by $\Gamma^{\prime}(+u_{j}^{\perp})$ we denote the graph induced in $\Gamma^{(n)}$ by the sets of vertices $U^{\prime}\cup\{u_{j}^{\perp}\}$ and $V^{\prime}\cup\{v^{\perp}_{j,i}:1\leq i\leq d-1\}$.
  • For any subgraph $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ of $\Gamma^{(n)}$ and any vertex $v\in V^{(n)}$ by $\Gamma^{\prime}(+v)$ (resp. $\Gamma^{\prime}(-v)$) we denote the graph induced in $\Gamma^{(n)}$ by the sets of vertices $U^{\prime}$ and $V^{\prime}\cup\{v\}$ (resp. $V^{\prime}\setminus\{v\}$).
  • By $h(x)$ we denote a witness of $c.e.H.h.c.(d)$ for $\Gamma$. At the end of each step, in Lemma 7.5 we produce $\widehat{h}(x)$ (possibly with some index) which is a witness of $c.e.H.h.c.(d)$ for $\Gamma^{(n)\star}$, defined with the help of $h$.
  • The elements adjacent to $u^{j+1}_{n}$ in the matching $\mathfrak{M}^{2}_{n}$ are denoted by $\dot{v}^{j+1}_{n,1},\ldots,\dot{v}^{j+1}_{n,d}$. The element $\dot{v}^{j+1}_{n,1}$ is a candidate for $v^{j+1}_{n,1}$.
  • $\mathcal{B}^{(n)}(u_{n})$ (resp. $\mathcal{S}^{(n)}(u_{n})$) is the ball (resp. sphere) in $\Gamma^{(n)}$ with the center $u_{n}$ and the radius $\max\{4h(3d(n+1))+3,5\}$.
  • $\mathcal{B}^{(n)}(u^{l+1}_{n})$ (resp. $\mathcal{S}^{(n)}(u^{l+1}_{n})$) is the ball (resp. sphere) in $\Gamma^{(n)}(-u^{0}_{n},\ldots,-u^{l}_{n})$ with the center $u^{l+1}_{n}$ and the radius $\max\{4h(3d(n+1))+3,5\}$.
  • $\mathfrak{M}^{1}_{n}$ denotes a finite $(1,d)$-matching in the bipartite graph $\mathcal{B}^{(n)\star}(u_{n})=\mathcal{B}^{(n)}(u_{n})\cap\Gamma^{(n)\star}$. It is obtained in such a way, that it satisfies the conditions of the perfect $(1,d)$-matchings for all vertices that are at the distance less than $\max\{4h(3d(n+1))+3,5\}$ from $u_{n}$.
  • $\mathfrak{M}^{2}_{n}$ denotes a finite $(1,d)$-matching in the bipartite graph $\mathcal{B}^{(n)\star}(u^{l+1}_{n})=\mathcal{B}^{(n)}(u^{l+1}_{n})\cap\Gamma^{(n)\star}$. It is obtained in such a way, that it satisfies the conditions of the perfect $(1,d)$-matchings for all vertices that are at the distance less than $\max\{4h(3d(n+1))+3,5\}$ from $u^{l+1}_{n}$.
  • The fan $(u^{\perp}_{n},\{v^{\perp}_{n,j}\,|1\leq j\leq d-1\})$ usually appears as a part of a fan of the matching $\mathfrak{M}^{2}_{n}$. We warn the reader that it is possible that $u^{\perp}_{n}$ does not exist.
  • We assume that all of these elements are natural numbers and are ordered according to the standard ordering of the natural numbers.
The main difference with the classical version of the construction from [Dud21] is that instead of $(1,d)$-matchings of infinite graphs we use $(1,d)$-matchings $\mathfrak{M}^{1}_{n}$, $\mathfrak{M}^{2}_{n}$ of finite subgraphs (when we produce $(1,d-1)$-fans to add to the constructed objects). In the Sections 6 and 7 we will show that the choice of the radius allows us to use this finite matchings $\mathfrak{M}^{1}_{n}$, $\mathfrak{M}^{2}_{n}$ without spoiling $c.e.H.h.c.(d)$ in the resulting graphs.
Before describing the construction in detail, we show the single step of the algorithm at the following scheme. We inform the reader that detailed illustrations of all parts of the step will be given at appropriate places.
5. Construction
From now on in this section we fix a highly computable bipartite graph $\Gamma=(U,V,E)$ which satisfies $c.e.H.h.c.(d)$ and the following conditions:
  • both $U$ and $V$ are identified with $\mathbb{N}\setminus\{0\}$;
  • $\Gamma$ is fully reflected;
  • $E$ does not contain edges of the form $(v,u_{v})$.
Before the construction of the proof of Theorem 4.1 we formulate an easy proposition which will be often used below without any additional comments. It gives an algorithm for finite matchings $\mathfrak{M}^{1}_{n}$ and $\mathfrak{M}^{2}_{n}$ in $\mathcal{B}^{(n)\star}(u^{0}_{n})$ and in $\mathcal{B}^{(n)\star}(u^{j}_{n})$ respectively (see the notation of Section 4.2).
Proposition 5.1.
Let $\Gamma^{\prime}=(U^{\prime},V^{\prime},E^{\prime})$ be a subgraph of $\Gamma=(U,V,E)$ which is also a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$. Then there is an uniform algorithm as follows.
Having an input $u^{*}\in U^{\prime}$ and an odd radius $r$ let $\mathcal{B}(u^{*})$ and $\mathcal{S}(u^{*})$ denote the corresponding ball and the sphere of radius $r$ in $\Gamma$. Put $\mathcal{B}^{\prime}(u^{*})=\mathcal{B}(u^{*})\cap\Gamma^{\prime}$ 111note that $\mathcal{B}^{\prime}(u^{*})$ does not have to be a ball in $\Gamma^{\prime}$ of the center $u^{*}$.
Then the algorithm finds a $(1,d)$-matching in $\mathcal{B}^{\prime}(u^{*})$, say $M_{u^{*}}$, which satisfies the conditions of perfect $(1,d)$-matchings for all $u\in U^{\prime}\cap\mathcal{B}(u^{*})$ and $v\in V^{\prime}\cap(\mathcal{B}(u^{*})\setminus\mathcal{S}(u^{*}))$.
From now on, such a matching $M_{u^{*}}$ will be called a perfect $(1,d)$-matching in the ball $\mathcal{B}^{\prime}(u^{*})$.
Proof.
All elements of $V^{\prime}\cap(\mathcal{B}(u^{*})\setminus\mathcal{S}(u^{*}))$ have the same neighbourhoods in $\Gamma^{\prime}$ and $\mathcal{B}^{\prime}(u^{*})$. By Theorem 3.5 there exists a perfect $(1,d)$-matching in $\Gamma^{\prime}$. Thus there exists a matching in $\mathcal{B}^{\prime}(u^{*})$ which satisfies the conditions of a perfect $(1,d)$-matchings for all $u\in U^{\prime}\cap\mathcal{B}(u^{*})$ and $v\in V^{\prime}\cap(\mathcal{B}(u^{*})\setminus\mathcal{S}(u^{*}))$. Since $\Gamma^{\prime}$ is highly computable and $\mathcal{B}^{\prime}(u^{*})$ is finite we find this matching in a computable way.
We now describe an inductive construction which is the heart of the proof of Theorem 4.1. Every step consists of three parts. Some of them are finished by claims stating that certain graphs satisfy Hall’s $d$-harem condition. These statements support further stages of the construction. We start with the detailed description of the first step of the construction.
5.1. Step 1, part 1
We take $u_{0}$, the first element of the set $U$, i.e. $u_{0}=1$. Let $\mathcal{B}^{(0)}(u_{0})$ be the ball centered at $u_{0}$ and of the radius $\max\{4h(3d)+3,5\}$. Apply Proposition 5.1 and compute a perfect $(1,d)$-matching $\mathfrak{M}^{1}_{0}$, in $\mathcal{B}^{(0)}(u_{0})$. Let $v^{0}_{0,1},\ldots,v^{0}_{0,d}$ be elements of $V$, ordered by its numbers, such that $(u_{0},v^{0}_{0,i})\in\mathfrak{M}^{1}_{0}$ for all $i\leq d$. Define the fan $M^{0}_{0}$ as the set of edges $(u_{0},v_{0,i})$ for $i\leq d-1$. Let
$\Gamma^{(0)}(-u_{0}):=(U\setminus\{u_{0}\},V\setminus\{v^{0}_{0,1},\ldots,v^{0}_{0,d-1}\}).$
Figure 1. The first part of step 1, the $\mathfrak{M}^{1}_{0}$-fan of $u_{0}$ where $M^{0}_{0}$ is green.
Claim 5.2.
The graph $\Gamma^{(0)}(-u_{0})$ satisfies $c.e.H.h.c.(d)$.
This claim follows from Lemma 7.1 which will be proved of in Section 7.
5.2. Remarks before part 2
Before the second part of step 1 let us discuss our local goals. Let $f_{0}$ be the partial function which corresponds to $M_{0}$ and let $\Gamma^{(1)}$ be the graph obtained after step 1, i.e. the part of $\Gamma$ after removal of $M_{0}$. We want to force that:
  1. (1)
    for all $n\in\mathsf{Dom}(f_{0})$ there exists $i$ such that $f_{0}^{i}(n)$ is a periodic point, and
  2. (2)
    $\Gamma^{(1)}$ is $U^{(1)}$-reflected.
It is clear that (1)-(2) are satisfied if we add the edge $(u_{v^{0}_{0,1}},v_{u_{0}})$ to the matching. This means that $f(v_{u_{0}})=u_{v^{0}_{0,1}}$, i.e. $f_{0}^{2}(u_{0})=f_{0}(f_{0}(v_{u_{0}}))=u_{0}$.
5.3. Step 1, part 2
Denote $u^{1}_{0}:=u_{v^{0}_{0,1}}$. Note that $(u^{1}_{0},v_{u_{0}})\in\Gamma^{(0)}(-u_{0})$, because $(u_{0},v^{0}_{0,1})$ is in $\Gamma^{(0)}$ and the latter one is $U^{(0)}$-reflected.
Remark 5.3.
For arbitrary $n$ part 2 of step $n$ depends on the subgraph $\Gamma^{(n-1)\perp}$. Since $\Gamma^{(0)\perp}$ is empty, at step 0 this part is reduced just to the choice of the vertex $u^{1}_{0}$.
In part 3 we will add the edge $(u^{1}_{0},v_{u_{0}})$, to $M_{0}^{1}$.
5.4. Step 1, part 3
Let $\mathcal{B}^{(0)}(u^{1}_{0})$ be the ball in $\Gamma^{(0)}(-u_{0})$ with the center $u^{1}_{0}$ and the radius $\max\{4h(3d)+3,5\}$. Since $\Gamma^{(0)}(-u_{0})$ satisfies $c.e.H.h.c.(d)$, we can apply Proposition 5.1 for $\Gamma^{(0)}(-u_{0})$ (viewed as $\Gamma=\Gamma^{\prime}$ in its formulation) and compute $\mathfrak{M}^{2}_{0}$, a perfect $(1,d)$-matching in $\mathcal{B}^{(0)}(u^{1}_{0})$. We remind the reader that $\dot{v}^{1}_{0,1},\ldots,\dot{v}^{1}_{1,d}$ are elements of $V$, ordered by their numbers, such that $(u^{1}_{0},\dot{v}^{1}_{0,i})\in\mathfrak{M}^{2}_{0}$ for all $i\leq d$.
Since $v_{u_{0}}$ is the least number in $V$, there are two possible cases:
  1. (1)
    $v_{u_{0}}=\dot{v}^{1}_{0,1}$. We set $v^{1}_{0,i}:=\dot{v}^{1}_{0,i}$, $11<i<d-$ 1 (i.e. $v^{1}_{0,1}=v_{u_{0}}$).
  2. (2)
    $v_{u_{0}}\neq\dot{v}^{1}_{0,1}$. Then find the fan $(u,v_{i})\in\mathfrak{M}^{2}_{0}$, $1\leq i\leq d$, such that $v_{u_{0}}=v_{1}$ (assuming that the ordering of $v_{i}$ corresponds to their indexes). We set:
    • $v^{1}_{0,1}:=v_{u_{0}}$ and $v^{1}_{0,i}:=\dot{v}^{1}_{0,i-1}$ for $12<i<d-$ 1,
    and define a candidate for $\Gamma^{(1)\perp}$:
    • $\dot{u}^{\perp}_{0}:=u$;
    • $\dot{v}^{\perp}_{0,i-1}:=v_{i}$ for $2<i<d$.
In either case define the fan $M^{1}_{0}$ as the set of edges $(u^{1}_{0},v^{1}_{0,i})$ for $11\leq i\leq d-$ 1.
Figure 2. Step 1, part 3. $\mathfrak{M}^{2}_{0}$ is red and $v_{u_{0}}$ is matched with $u$. We want the edge $(u^{1}_{0},v_{u_{0}})$ to be in $M_{0}$. Force the situation from Figure 3.
Figure 3. $M^{1}_{0}$ is green. It is possible that the purple fan consisting of edges $(\dot{u}^{\perp}_{0},\dot{v}^{\perp}_{0,1}),(\dot{u}^{\perp}_{0},\dot{v}^{\perp}_{0,2})$ will be added to $\Gamma^{(1)\perp}$.
Put $M_{0}=M^{0}_{0}\cup M^{1}_{0}$. We obtain $\Gamma^{(1)}$ by removal of $M_{0}$ from $\Gamma$. Since $u_{0},v_{u_{0}},v^{0}_{0,1}$ and $u^{1}_{0}=u_{v^{0}_{0,1}}$ have been removed, $\Gamma^{(1)}$ is $U^{(1)}$-reflected. It might turn out that it does not satisfy Hall’s $d$-harem condition.
Let $\Gamma^{\prime}_{0}=(U^{(1)}\setminus\{\dot{u}_{0}^{\perp}\},V^{(1)}\setminus\{\dot{v}^{\perp}_{0,1},...,\dot{v}^{\perp}_{0,d-1}\})$. The statement below follows from Lemma 7.5.
Claim 5.4.
At least one of $\Gamma^{\prime}_{0}$ or $\Gamma^{(1)}$ satisfies $c.e.H.h.c.(d)$.
5.5. Output of the first step
If $\Gamma^{(1)}$ satisfies $c.e.H.h.c.(d)$, set $\Gamma^{(1)\star}:=\Gamma^{(1)}$ and $U^{(1)\perp}=\emptyset,V^{(1)\perp}=\emptyset$. If $\Gamma^{(1)}$ does not satisfy $c.e.H.h.c.(d)$, set $\Gamma^{(1)\star}:=\Gamma^{\prime}_{0}$ and
  • $u^{\perp}_{0}:=\dot{u}^{\perp}_{0}$;
  • $v^{\perp}_{0,i}:=\dot{v}^{\perp}_{0,i}$;
  • $U^{(1)\perp}=\{u^{\perp}_{0}\}$;
  • $V^{(1)\perp}=\{v^{\perp}_{0,i}:1\leq i\leq d-1\}.$
5.6. Remarks before step $n+1$
Before description of the $n+1$-st step let us discuss the situation after the $n$-th step.
We have constructed graphs $\Gamma^{(n)}$ and $\Gamma^{(n)\star}$, where $\Gamma^{(n)}$ is $U^{(n)}$-reflected and $\Gamma^{(n)\star}$ satisfies $c.e.H.h.c.(d)$. Also note that since $|U^{(n)\perp}\setminus U^{(n-1)\perp}|\leq 1$, there are at most $n$ roots $u^{\perp}_{i}$ of fans in $U^{(n)\perp}$ (see Section 2.4 for the corresponding definition).
5.7. Step n+1, part 1
Take $u_{n}$, the first element of the set $U^{(n)}$. In order to define $M^{0}_{n}$ we have two possible cases:
  1. (1)
    There is $j$ such that $u_{n}=u_{j}^{\perp}\in U^{(n)\perp}$. Then we set $M^{0}_{n}$ to consist of all edges of the form $(u^{\perp}_{j},v^{\perp}_{j,i})$ and remove the fan with the root $u_{j}^{\perp}$ from $\Gamma^{(n)\perp}$. We redefine $U^{(n)\perp}$ and $V^{(n)\perp}$ accordingly (in particular $u_{j}^{\perp}$ is removed from $U^{(n)\perp}$).
  2. (2)
    If $u_{n}\not\in U^{(n)\perp}$, then take the ball $\mathcal{B}^{(n)}(u_{n})$ in $\Gamma^{(n)}$ with the center $u_{n}$ and the radius $\max\{4h(d(3n+1))+3,5\}$. Applying Proposition 5.1 compute $\mathfrak{M}^{1}_{n}$, a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u_{n}):=\mathcal{B}^{(n)}(u_{n})\cap\Gamma^{(n)\star}$. Verify whether there is $j$ with $(u_{v^{0}_{n,j}},v_{u_{n}})\in\Gamma^{(n)\perp}$. By the definition of $\Gamma^{(n)\perp}$ it can happen for at most one $j$ [^2] Note here that it can also happen that $v_{u_{n}}$ is not even in $\Gamma^{(n)}$. If there is such $j$ then it can be equal to $d$ or not. In the first case set $M^{0}_{n}$ to consist of edges $(u_{n},v^{0}_{n,i})\in\mathfrak{M}^{1}_{n}$ for $i\neq d-1$. In all other possibilities we set $M^{0}_{n}$ to consist of edges $(u_{n},v^{0}_{n,i})\in\mathfrak{M}^{1}_{n}$ for $i\leq d-1$.
Remark 5.5.
Note that as a result in the case of existence of $j$ as above the edge $(u_{n},v^{0}_{n,j})$ is included into $M^{0}_{n}$. Although the final part of the algorithm of (2) can be presented easier, see Figure 4, we do not want to miss this point.
Figure 4. Detailed version of the algorithm used at part 1 of step $n+1$.
The following claim follows from Lemma 7.1 below.
Claim 5.6.
Let $\Gamma^{(n)\star}(-u_{n})$ be $\Gamma^{(n)}(-u_{n})\cap\Gamma^{(n)\star}$. One of the following holds:
  • $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$;
  • we can compute $u_{j}^{\perp}\in U^{(n)\perp}$ such that the graph $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$.
5.8. Output of part 1
This is $M^{0}_{n}$. We also update our graphs in the following situation. If $\Gamma^{(n)\star}(-u_{n})$ does not satisfy $c.e.H.h.c.(d)$, let $u_{j}^{\perp}$ be an element from $U^{(n)\perp}$ realizing the second possibility of Claim 5.6. We remove the fan of $u^{\perp}_{j}$ with its leaves from $(U^{(n)\perp},V^{(n)\perp})$ and then we put it into $\Gamma^{(n)\star}$. Thus the latter graph (and $U^{(n)\perp},V^{(n)\perp}$) are updated. It is clear that now the redefined $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$.
5.9. Remarks before part 2
Before the rest of the step $n+1$, we describe the goals which we want to achieve after the step:
  1. (1)
    the partial function $f_{n}$ corresponding to $\bigcup\limits_{i=0}^{n}M_{i}$ has controlled sizes of its cycles, and
  2. (2)
    the graph $\Gamma^{(n+1)}$ obtained at the end of the step is $U^{(n+1)}$-reflected and the corresponding graph $\Gamma^{(n+1)\star}$ satisfies Hall’s $d$-harem condition.
In order to achieve the first condition we will organize one of the the following properties:
  1. (i)
    there is a sequence of vertices $u^{0}_{n},u^{1}_{n},u^{2}_{n},\ldots,u^{j}_{n}$, $1\leq j\leq n$ such that each edge $(u^{i}_{n},v_{u^{i-1}_{n}})$ belongs to $M_{n}$, and for some $10\leq\ell\leq j-$ 1 the sequence $(u^{\ell}_{n},u^{\ell+1}_{n},\ldots,u^{j}_{n})$ is a cycle;
  2. (ii)
    there is a sequence of vertices $u^{0}_{n},u^{1}_{n},u^{2}_{n},\ldots,u^{j}_{n}$, $j\leq n$ such that each edge $(u^{i}_{n},v_{u^{i-1}_{n}})$ belongs to $M_{n}$, and $v_{u^{j}_{n}}$ is already adjacent to some edge from $\bigcup\limits_{i=0}^{n-1}M_{i}$;
5.10. Step n+1, part 2
We begin by checking whether $v_{u_{n}}$ belongs to $\Gamma^{(n)\perp}$. If $v_{u_{n}}\in\Gamma^{(n)\perp}$ then we denote $u_{n}$ by $u^{0}_{n}$ and begin the following process of choosing the consecutive vertices $u^{i}_{n}$.
First step of iteration. Assume that for some $j_{0},i$ we have $v_{u_{n}}=v^{\perp}_{j_{0},i}\in V^{(n)\perp}$. Then we set
$u^{1}_{n}=u^{\perp}_{j_{0}}$, $v^{1}_{n,k}:=v^{\perp}_{j_{0},k}$, $k\leq d-1$,
$M^{1}_{n}:=\{(u^{1}_{n},v^{1}_{n,k}):1\leq k\leq d-1\}$
and check whether $v_{u^{1}_{n}}\in\Gamma^{(n)\perp}(-u^{0}_{n})$.
If it is so we repeat the iteration for $v_{u^{1}_{n}}$. Note that $v_{u^{1}_{n}}\in\Gamma^{(n)\perp}(-u^{0}_{n},-u^{1}_{n})$ then.
Single step of iteration. We verify if $v_{u^{m}_{n}}\in\Gamma^{(n)\perp}(-u^{0}_{n},-u^{1}_{n},\ldots,-u^{m}_{n})$. If it is so then for some $j_{m},i$ we have $v_{u^{m}_{n}}=v^{\perp}_{j_{m},i}\in V^{(n)\perp}$. Define
$u^{m+1}_{n}=u^{\perp}_{j_{m}}$, $v^{m+1}_{n,k}:=v^{\perp}_{j_{m},k}$, $11\leq k\leq d-$ 1,
$M^{m}_{n}:=\{(u^{m+1}_{n},v^{m+1}_{n,k}):1\leq k\leq d-1\}$
and we repeat the iteration for $v_{u^{m+1}_{n}}$. This ends the single iteration step.
Since $|U^{(n)\perp}|\leq n$, the procedure ends after at most $n$ iterations. Therefore one of the following cases is realized for some $l\leq n$:
  1. (1)
    $v_{u^{l}_{n}}\notin\Gamma^{(n)}$;
  2. (2)
    $v_{u^{l}_{n}}\in\Gamma^{(n)}$, but $v_{u^{l}_{n}}\notin\Gamma^{(n)}(-u^{0}_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$ (this case is impossible for $l=0$);
  3. (3)
    $v_{u^{l}_{n}}\in\Gamma^{(n)\star}$.
In case (1) $v_{u^{l}_{n}}$ was already added to $M$ at preceding steps and condition (ii) of Remarks at point 5.9 is satisfied. We finish part 2 of step $n+1$ without a new cycle.
In case (2) the last iteration closes the cycle $(u^{k}_{n},\ldots,u^{l}_{n})$. The length of this cycle is not greater than $l+1$.
In each of these two cases we skip part 3 of the step and go to the output of step $n+1$, see 5.13. We set $M_{n}=\bigcup\limits_{k=0}^{l}M^{k}_{n}$ and obtain the graph $\Gamma^{(n+1)}$ from $\Gamma^{(n)}$ by removal of $M_{n}$-fans. Since for every $u\in U^{(n)}\setminus U^{(n+1)}$ the element $v_{u}$ is either removed as well or was not in $V^{(n)}$ from the beginning, then the graph is $U^{(n+1)}$-reflected.
Case (3) is the most complicated case. We finish this part of the step by taking a new term $u^{l+1}_{n}:=u_{v^{l}_{n,1}}$. We force a cycle $(u^{l}_{n},u^{l+1}_{n})$ of length $2$ in part 3 of the step.
Figure 5 illustrates the algorithm of the second part of the step.
Figure 5. Detailed version of the algorithm of part 2 of step $n+1$.
5.11. Example
Before we move to the third part of the step, we give an example of a cycle obtained by construction in case (2) of this part of the step.
The following pictures show how a cycle of length $3$ can be obtained by this procedure in the graph satisfying Hall’s $3$-harem condition.
Figure 6. $\Gamma^{(n)\star}$ is black, $\Gamma^{(n)\perp}$ is purple.
Figure 7. $\mathfrak{M}^{n}_{1}$ is red and green, $M_{n}^{0}$ is green. We have $v_{u_{n}}=v^{\perp}_{j_{0},2}$.
Figure 8. $M_{n}^{1}$ is green. We have $v_{u^{\perp}_{j_{1}}}=v^{\perp}_{j_{2},2}$. Moreover $v_{u_{j_{1}}^{\perp}}=v^{0}_{n,2}$.
To demonstrate the cycle in a single picture, we take 4 copies of $\mathbb{N}$ as rows. The first and the second ones correspond to $V$ and $U$ after part 1 of the step. The second and the third rows correspond to the sets $V$ and $U$ after the first iteration step. The third and the fourth rows correspond to $U$ and $V$ after the second step of the iteration.
Figure 9. The final $M_{n}$ with the obtained cycle of length $3$ marked in blue. The red edge is the edge from $\mathfrak{M}^{n}_{1}$ that was not added to $M_{n}^{0}$.
5.12. Step n+1, part 3
Let $\mathcal{B}^{(n)}(u^{l+1}_{n})$ be the ball in $\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$ with the center $u^{l+1}_{n}$ and the radius $\max\{4h(3d(n+1))+3,5\}$. Let $\mathcal{B}^{(n)\star}(u^{l+1}_{n}):=\mathcal{B}^{(n)}(u^{l+1}_{n})\cap\Gamma^{(n)\star}$.
In the last case the edge $(u^{l}_{n},v^{l}_{n,1})$ is in $\Gamma^{(n)}$ and $v_{u^{l}_{n}}$ is not in $\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$, thus applying $U^{(n)}$-reflectedness of $\Gamma^{(n)}$ we see $(u^{l+1}_{n},v_{u^{l}_{n}})\in\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$. Observe that since $u^{\perp}_{j_{1}},\ldots,u^{\perp}_{j_{l}}$ are in $U^{(n)\perp}$,
$\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})\cap\Gamma^{(n)\star}=\Gamma^{(n)}(-u_{n})\cap\Gamma^{(n)\star}=\Gamma^{(n)\star}(-u_{n}).$
By part 1 of this step the graph $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$, thus we can apply Proposition 5.1 for $\Gamma^{(n)\star}(-u_{n})$ (viewed as $\Gamma^{\prime}$ in formulation, with $\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$ viewed as $\Gamma$) and compute $\mathfrak{M}^{2}_{n}$, a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u^{l+1}_{n})$.
We now check whether there is $j$ with $u^{l+1}_{n}=u^{\perp}_{j}$. If it exists, we set $\dot{v}^{l+1}_{n,i}:=v^{\perp}_{j,i}$ for $11\leq i\leq d-$ 1. When it does not exist, $\dot{v}^{l+1}_{n,1},\ldots,\dot{v}^{l+1}_{n,d}$ will denote the elements adjacent to $u^{l+1}_{n}$ under $\mathfrak{M}^{2}_{n}$.
There are two cases:
  1. A)
    $(u^{l+1}_{n},v_{u^{l}_{n}})\in\mathfrak{M}^{2}_{n}$, i.e. $v_{u^{l}_{n}}=\dot{v}^{l+1}_{n,k}$ for some $1\leq k\leq d$. In this case $u^{l+1}_{n}$ cannot be $u^{\perp}_{j}$ for any $j$.
  2. B)
    $(u^{l+1}_{n},v_{u^{l}_{n}})\notin\mathfrak{M}^{2}_{n}$, i.e. there exists some $u\in\Gamma^{(n)\star}(-u_{n})$, such that $v_{u^{l}_{n}}=v_{k}$ for some $1\leq k\leq d$, where $v_{1},\ldots,v_{d}$ denote the elements adjacent to $u$ under $\mathfrak{M}^{2}_{n}$. In this case it is possible that $u^{l+1}_{n}$ coincides with some $u^{\perp}_{j}$.
In either case we produce a cycle of length 2 by including the pair $(u^{l+1}_{n},v_{u^{l}_{n}})$ into $M^{l+1}_{n}$. In fact we include it into $M^{l+1}_{n}$ together with a fan with the root $u^{l+1}_{n}$ and $(d-2)$ leaves taken among $\dot{v}^{l+1}_{n,i}$.
To be precise we organize it as follows.
In case A) If $k=d$, we set $v^{l+1}_{n,i-1}:=\dot{v}^{l+1}_{n,i}$ for $2\leq i\leq d$. If $k\neq d$ we set $v^{l+1}_{n,i}:=\dot{v}^{l+1}_{n,i}$ for $11\leq i\leq d-$ 1. The set $M^{l+1}_{n}$ consists of edges $(u^{l+1}_{n},v^{l+1}_{n,i})$ for $11\leq i\leq d-$ 1. The procedure is finished.
In case B) $v^{l+1}_{n,1}=v_{u^{l}_{n}}$ and $v^{l+1}_{n,i+1}=\dot{v}^{l+1}_{n,i}$, $21\leq i\leq d-$ 2.333It is worth noting here that if $u^{l+1}_{n}$ coincides with some $u^{\perp}_{j}$, then the vertex $\dot{v}^{l+1}_{n,d}$ does not exist, i.e the only vertex from the fan of $u^{l+1}_{n}$ that is outside of $M^{l+1}_{n}$ is $\dot{v}^{l+1}_{n,d-1}$. We define $\dot{u}_{n}^{\perp}:=u$ and rename the remaining $d-1$ vertices $v_{j}$ to $\dot{v}^{\perp}_{n,i}$. The procedure is finished.
Figure 10 illustrates part 3 of the algorithm.
Figure 10. Detailed version of the algorithm used at the part 3 of the step $n+1$.
Let $M_{n}=\bigcup\limits_{k=0}^{l+1}M^{k}_{n}$. We obtain the graph $\Gamma^{(n+1)}$ from $\Gamma^{(n)}$ by removal of $M_{n}$-fans. Since for each $u\in U^{(n)}\setminus U^{(n+1)}$ the element $v_{u}$ is also removed, then the graph is $U^{(n+1)}$-reflected.
5.13. Output of the step
This is $M_{n}$ and $\Gamma^{(n+1)}$. It remains to produce graphs $\Gamma^{(n+1)\star},\Gamma^{(n+1)\perp}$. To do that we define auxilliary graphs $\mathfrak{T}$ and $\dot{\Gamma}^{(n)\perp}$.
In cases (1), (2) and (3)A) we set $\mathfrak{U}:=U^{(n+1)}\setminus U^{(n)\perp},\mathfrak{V}:=V^{(n+1)}\setminus V^{(n)\perp}$. Let $\mathfrak{T}=(\mathfrak{U},\mathfrak{V})$ and $\dot{\Gamma}^{(n)\perp}:=\Gamma^{(n)\perp}\cap\Gamma^{(n+1)}$. Note here that in cases (1) and (2) only elements of $\Gamma^{(n)\perp}$ were added to $M_{n}$ at the part 2 of the step. Therefore in these cases $\mathfrak{T}$ coincides with $\Gamma^{(n)\star}(-u_{n})$ and satisfies Hall’s $d$-harem condition.
In case (3)B) we set $\mathfrak{U}:=U^{(n+1)}\setminus(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\}),\mathfrak{V}:=V^{(n+1)}\setminus(V^{(n)\perp}\cup\{\dot{v}_{n,i}^{\perp}:1\leq i\leq d-1\})$. Let $\mathfrak{T}=(\mathfrak{U},\mathfrak{V})$ and $\dot{U}^{(n)\perp}:=(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\})\cap U^{(n+1)}$. We define $\dot{V}^{(n)\perp}$ accordingly: $\dot{V}^{(n)\perp}:=(V^{(n)\perp}\cup\{\dot{v}_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}$, and set
$\dot{\Gamma}^{(n)\perp}:=(\dot{U}^{(n)\perp},\dot{V}^{(n)\perp}).$
The following claim follows from Lemma 7.5 below.
Claim 5.7.
At least one of the following holds:
  • $\mathfrak{T}$ satisfies $c.e.H.h.c.(d)$;
  • there exists some vertex $u_{j}^{\perp}\in\dot{U}^{(n)\perp}$ such that the graph $\mathfrak{T}(+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$.
  • there exist vertices $u_{i}^{\perp},u_{j}^{\perp}\in\dot{U}^{(n)\perp}$ such that the graph $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$.
Now, depending on the output of the claim above we define graphs $\Gamma^{(n+1)\star},\Gamma^{(n+1)\perp}$.
In the case (3)B) (i.e. $\dot{u}_{n}^{\perp}$ exists) if $u_{i}^{\perp}\neq\dot{u}_{n}^{\perp}\neq u_{j}^{\perp}$ (for any output of the claim), we set $u_{n}^{\perp}:=\dot{u}_{n}^{\perp}$ and $v_{n,k}^{\perp}:=\dot{v}_{n,k}^{\perp}$ for $11\leq k\leq d-$ 1. Otherwise, or in the remaining cases $u_{n}^{\perp},v_{n,k}^{\perp}$ are not defined.
In the first case of the claim we set $\Gamma^{(n+1)\star}:=\mathfrak{T}$ and
$U^{(n+1)\perp}:=(U^{(n)\perp}\cup\{u_{n}^{\perp}\})\cap U^{(n+1)},$
$V^{(n+1)\perp}:=(V^{(n)\perp}\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
In the second case we set $\Gamma^{(n+1)\star}:=\mathfrak{T}(+u_{j}^{\perp})$and
$U^{(n+1)\perp}:=((U^{(n)\perp}\setminus\{u_{j}^{\perp}\})\cup\{u_{n}^{\perp}\})\cap U^{(n+1)},$
$V^{(n+1)\perp}:=((V^{(n)\perp}\setminus\{v_{j,k}^{\perp}:1\leq k\leq d-1\})\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
In the third case we set $\Gamma^{(n+1)\star}:=\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ and
$U^{(n+1)\perp}:=((U^{(n)\perp}\setminus\{u_{i}^{\perp},u_{j}^{\perp}\})\cup\{u_{n}^{\perp}\})\cap U^{(n+1)},$
$V^{(n+1)\perp}:=((V^{(n)\perp}\setminus\{v_{i,k}^{\perp},v_{j,k}^{\perp}:1\leq k\leq d-1\})\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
6. Technical Lemmas
The notation used in this section does not correspond to the construction above.
Throughout this section $\Gamma=(U,V,E)$ denotes a bipartite graph
and $\Gamma^{\perp}=(U^{\perp},V^{\perp},E^{\perp})$ denotes its subgraph.
The graph $\Gamma^{\star}=(U^{\star},V^{\star},E^{\star})$ is an induced subgraph of $\Gamma$ such that
$U^{\star}\cap U^{\perp}=\emptyset=V^{\star}\cap V^{\perp}.$
Below we always assume that $d$ is a natural number greater than $1$.
The following situation will arise several times in our arguments in Section 7. Let $X$ be a subset of $V$ such that
$|N_{\Gamma}(X)|\geq\frac{1}{d}|X|,$
but
$|N_{\Gamma}(X)\setminus U^{\perp}|<\frac{1}{d}|X|.$
Thus we can conclude that $N_{\Gamma}(X)\cap U^{\perp}\neq\emptyset$.
The following lemma describes typical circumstances which lead to this situation.
Lemma 6.1.
Let $\Gamma=(U,V,E)$ be $U$-reflected and let $\Gamma^{\star}$ be a subgraph of $\Gamma$ induced by the sets of vertices $U^{\star}=U\setminus U^{\perp}$, $V^{\star}=V\setminus V^{\perp}$. Assume that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition and assume that for each $Y\subset U^{\perp}$ we have $|N_{\Gamma^{\perp}}(Y)|\geq(d-1)|Y|$.
Then for any $X\subseteq V$ we have $|N_{\Gamma}(X)|\geq(d-1)|X|$. In particular $|N_{\Gamma}(X)|\geq\frac{1}{d}|X|$
Proof.
Let $U_{X}:=\{u\in U:v_{u}\in X\}$. By $U$-reflectedness of $\Gamma$ we have $|U_{X}|=|X|$. Consider the sets
$U_{X}^{\star}:=U_{X}\setminus U^{\perp}$
and
$U_{X}^{\perp}:=U_{X}\cap U^{\perp}.$
Using $U$-reflectedness of $\Gamma$ again, we see
$|N_{\Gamma}(X)|\geq|N_{\Gamma^{\star}}(U_{X}^{\star})|+|N_{\Gamma}(U^{\perp}_{X})\cap V^{\perp}|.$
Since Hall’s $d$-harem condition is satisfied for any subset of $U^{\star}$,
$|N_{\Gamma^{\star}}(U_{X}^{\star})|\geq d|U_{X}^{\star}|.$
Moreover we have $|N_{\Gamma}(U_{X}^{\perp})\cap V^{\perp}|\geq(d-1)|U_{X}^{\perp}|$. Therefore
$|N_{\Gamma^{\star}}(U_{X}^{\star})|+|N_{\Gamma}(U^{\perp}_{X})\cap V^{\perp}|\geq(d-1)|U_{X}|.$
Since $d\geq 2$ it follows that
$|N_{\Gamma}(X)|\geq(d-1)|X|\geq\frac{1}{d}|X|.$
This lemma cannot be applied in the situation of Part 1 of the step of the main construction, since some vertices from a $U$-reflected graph are removed. In that case, we have a graph $\Gamma$ which is not $U$-reflected, but for some $v\in V$ its subgraph $\Gamma(-v)$ is $U$-reflected. Because of this, in the following lemmas both $\Gamma^{\perp}$ and $\Gamma^{\star}$ are the subgraphs of $\Gamma(-v)$.
Lemma 6.2.
Let $v$ be a vertex from $V$ and $\Gamma(-v)=(U,V(-v))$. Assume that $\Gamma(-v)$ is $U$-reflected. Let $\Gamma^{\perp}=(U^{\perp},V^{\perp},E^{\perp})$ be a subgraph of $\Gamma(-v)$ and let $\Gamma^{\star}$ be defined as the subgraph of $\Gamma(-v)$ induced by the sets of vertices $U^{\star}:=U\setminus U^{\perp},V^{\star}:=V(-v)\setminus V^{\perp}$.
Assume that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition and assume that for each $Y\subset U^{\perp}$ we have $|N_{\Gamma}(Y)\cap V^{\perp}|\geq(d-1)|Y|$.
Then for any $X\subseteq V\setminus\{v\}$ the inequality $|N_{\Gamma}(X)|\geq(d-1)|X|-1$ holds.
Moreover, the equality $|N_{\Gamma}(X)|=(d-1)|X|-1$ can happen only if $v\in N_{\Gamma}(U_{X})$ and $N_{\Gamma^{\star}}(U_{X})=\emptyset$, where $U_{X}:=\{u\in U:v_{u}\in X\}$.
Proof.
By $U$-reflectedness of $\Gamma(-v)$ we have $|U_{X}|=|X|$ and
$|N_{\Gamma}(X)|+1\geq|N_{\Gamma(-v)}(U_{X})|.$
Furthermore, if $v\notin N_{\Gamma}(U_{X})$, then
$|N_{\Gamma}(X)|\geq|N_{\Gamma(-v)}(U_{X})|.$
Consider the sets
$U_{X}^{\star}:=U_{X}\setminus U^{\perp}\,\mbox{ and }\,U_{X}^{\perp}:=U_{X}\cap U^{\perp}.$
Applying the argument of Lemma 6.1 to $\Gamma(-v)$ we obtain
$|N_{\Gamma(-v)}(U_{X})|\geq|N_{\Gamma^{\star}}(U_{X}^{\star})|+|N_{\Gamma}(U^{\perp}_{X})\cap V^{\perp}|\geq(d-1)|U_{X}|.$
Thus $|N_{\Gamma}(X)|\geq|N_{\Gamma(-v)}(U_{X})|-1\geq(d-1)|X|-1$.
Observe that $|N_{\Gamma(-v)}(U_{X})|=(d-1)|U_{X}|$ only if $N_{\Gamma(-v)}(U_{X})\subseteq V^{\perp}$, i.e. $N_{\Gamma^{\star}}(U_{X})=\emptyset$. If $v\notin N_{\Gamma}(U_{X})$ then
$|N_{\Gamma}(X)|\geq|N_{\Gamma(-v)}(U_{X})|\geq(d-1)|X|.$
Therefore $|N_{\Gamma}(X)|=(d-1)|X|-1$ only if $v\in N_{\Gamma}(U_{X})$ and $N_{\Gamma^{\star}}(U_{X})=\emptyset$.
Another fact that will be regularly used in the proofs of Lemmas 7.1 and 7.5 is the following.
Proposition 6.3 (Removing at most $k$ fans from a ball).
Assume that $\Gamma=(U,V,E)$ is a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$ with witness $h$. Let
  • $u\in U$ and $\mathcal{B}(u)$ be the ball in $\Gamma$ with the center $u$ and the radius of at least
    $\max\{2h(kd)+3,5\}$;
  • $M_{u}$ be a perfect $(1,d)$-matching in the ball $\mathcal{B}(u)$;
  • $U_{1}$ be a subset of $U$ from the ball of $u$ of radius $\leq 2$, $|U_{1}|\leq k$, and $V_{1}$ be the set of vertices adjacent in $M_{u}$ to vertices from $U_{1}$.
Then $\Gamma^{\prime}:=(U\setminus U_{1},V\setminus V_{1})$ satisfies Hall’s $d$-harem condition.
Proof.
Let us show that the inequality $|N_{\Gamma^{\prime}}(X)|-d|X|\geq 0$ holds for any $X\subset U\setminus U_{1}$. If $X$ is not connected, then its neighbourhood is a disjoint union of the neighbourhoods of the connected components of $X$. Thus the inequality for each component implies it for $X$. Therefore without loss of generality we may also assume that $X$ is connected. Further, we may assume that $X\nsubseteq U\cap\mathcal{B}(u)$ and $V_{1}\cap N_{\Gamma}(X)\neq\emptyset$. Indeed, if $V_{1}\cap N_{\Gamma}(X)=\emptyset$ then $N_{\Gamma^{\prime}}(X)=N_{\Gamma}(X)$, i.e. $|N_{\Gamma^{\prime}}(X)|\geq d|X|$. On the other hand if $X\subset U\cap\mathcal{B}(u)$ then using existence of an $(1,d)$-matching $M_{u}$ from $U\cap\mathcal{B}(u)$ to $V\cap\mathcal{B}(u)$ which is perfect for subsets of $U\cap\mathcal{B}(u)$, we see that $|N_{\Gamma^{\prime}}(X)|-d|X|\geq 0$.
Let us choose $u_{1},u_{2}\in X$ such that $u_{1}\in N_{\Gamma}(V_{1})$ and $u_{2}\in X\setminus\mathcal{B}(u)$. By the choice of the radius of the ball $\mathcal{B}(u)$ the distance between $u_{1}$ and $u_{2}$ is at least $12h(2d)+$ 1. Thus $|X|\geq h(kd)$. Since $\Gamma$ satisfies $c.e.H.h.c.(d)$ we have
$|N_{\Gamma}(X)|-d|X|\geq kd.$
We know that $|N_{\Gamma^{\prime}}(X)|\geq|N_{\Gamma}(X)|-kd$, therefore
$|N_{\Gamma^{\prime}}(X)|-d|X|\geq|N_{\Gamma}(X)|-kd-d|X|\geq 0.$
The proof for the case of subsets of $V^{\prime}$ is completely analogous.
The following definition will be useful in the proofs of the lemmas of the next section.
Definition 6.4.
Let $\Gamma:=(U,V,E)$ be a bipartite graph, and $\Gamma^{\star}=(U^{\star},V^{\star},E^{\star})$ be a subgraph in $\Gamma$ satisfying Hall’s $d$-harem condition. Assume that $M$ is a corresponding perfect $(1,d)$-matching in $\Gamma^{\star}$.
[…]
  • $v^{\prime}_{n}=x$;
  • $(u^{\prime}_{i},v^{\prime}_{i})\in M$ and $(u^{\prime}_{i},v^{\prime}_{i+1})\in E\setminus M$, where $i<n$;
  • $(y,v^{\prime}_{0})\in E$.
In the technical lemma below we use the following notion of Kierstead, [Kie83]. The subset $X$ of $U$ (resp. of $V$) is called connected if for all $x,x^{\prime}\in X$ there exist a path $x=p_{0},p_{1},\ldots,p_{k}=x^{\prime}$ in $\Gamma$ such that $p_{i}\in X\cup N_{\Gamma}(X)$ for all $i\leq k$. Note here that this definition concerns only bipartite graphs.
Lemma 6.5.
Let $\Gamma=(U,V,E)$ be a bipartite graph. Let $\Gamma^{\star}=(U^{\star},V^{\star},E^{\star})$ be a subgraph of $\Gamma$ satisfying Hall’s $d$-harem condition and $M$ be a corresponding perfect $(1,d)$-matching.
Assume that
  • $\widehat{v}\in N_{\Gamma}(U^{\star})\setminus V^{\star}$;
  • $X$ is a minimal connected subset in $V^{\star}\cup\{\widehat{v}\}$ with $|N_{\Gamma}(X)|<\frac{1}{d}|X|$;
  • $\widehat{u}\in N_{\Gamma}(X)\setminus U^{\star}$;
[…]
Proof.
It is clear that $\widehat{v}\in X$. Let $X^{\prime}$ denote the subset of all elements of $X$ that are accessible from $\widehat{u}$ through $X$ by the matching $M$. Assume that $\widehat{v}\not\in X^{\prime}$. Since $\widehat{u}\in N_{\Gamma}(X)$, we then see that there exist $u^{\prime}_{0},v^{\prime}_{0}$ such that $\widehat{u},u^{\prime}_{0}\in N_{\Gamma}(v^{\prime}_{0})$ and $(u^{\prime}_{0},v^{\prime}_{0})\in M$. Therefore $X^{\prime}\neq\emptyset$. We will show that $|N_{\Gamma}(X\setminus X^{\prime})|<\frac{1}{d}|X\setminus X^{\prime}|$.
Let $|X^{\prime}|=l$ and let $U_{M}(X^{\prime})=\{u\in U\,|\,(\exists v\in X^{\prime})(u,v)\in M\}$. Since elements of $X\setminus X^{\prime}$ are not accessible from $\widehat{u}$ through $X$ by the matching $M$, then $N_{\Gamma}(X\setminus X^{\prime})\cap U_{M}(X^{\prime})=\emptyset$.
Using this we see that
$|N_{\Gamma}(X\setminus X^{\prime})|\leq|N_{\Gamma}(X)|-|U_{M}(X^{\prime})|.$
Since $M$ is a $(1,d)$-matching, each element of $U_{M}(X^{\prime})$ can be matched with at most $d$ elements from $X^{\prime}$. Therefore $|U_{M}(X^{\prime})|\geq\lceil\frac{l}{d}\rceil$ and
$|N_{\Gamma}(X\setminus X^{\prime})|\leq N_{\Gamma}(X)|-|U_{M}(X^{\prime})|\leq|N_{\Gamma}(X)|-\lceil\frac{l}{d}\rceil<\frac{1}{d}(|X|-l)=\frac{1}{d}|X\setminus X^{\prime}|.$
[…]
This lemma will be essential in the proofs below in order to replace some $(1,d-1)$-fans in the matching. Since in our case this matching is partial and finite we need an additional technical lemma. It will guarantee that in a graph satisfying $c.e.H.h.c.(d)$ Hall’s $d$-harem condition can be preserved after some replacement of $(1,d-1)$-fans.
Lemma 6.6 (Replacement of a fan in a finite matching).
Let
  • $\Gamma=(U,V,E)$ be a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$ with a witness $h$;
  • $\Gamma^{\star}=(U^{\star},V^{\star})$ be a subgraph of $\Gamma$ such that $\Gamma\setminus\Gamma^{\star}$ consists of $n$ $(1,d-1)$-fans;
  • $u\in U$ and $\mathcal{B}(u)$ be a ball in $\Gamma$ of $u$ of odd radius of at least $\max\{4h(d(n+1))+3,5\}$;
  • $\mathfrak{F}_{u}$ be an $(1,d-1)$-fan in $\Gamma^{\star}$ with the root $u$;
  • $\mathfrak{S}$ be an $(1,d-1)$-fan in $\Gamma\setminus\Gamma^{\star}$ with the root $u^{\perp}$;
  • $\mathfrak{M}$ be a perfect $(1,d)$-matching in the ball $((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})\cap\mathcal{B}(u)$.
If $\Gamma^{\star}$ satisfies $c.e.H.h.c.(d)$ with the witness
$\widehat{h}(m)=\left\{\begin{array}[]{rr}0,\quad\text{if}\quad m=0,\\ h(m+nd),\quad\text{if}\quad m>0.\end{array}\right.$
but $\Gamma^{\star}\setminus\mathfrak{F}_{u}$ does not satisfy Hall’s $d$-harem condition, then $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ satisfies Hall’s $d$-harem condition.
Proof.
We start the proof with several claims.
Claim 6.7.
All subsets of $U^{\star}\setminus\{u\}$ satisfy Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$.
Proof of Claim 6.7.
Assume that $X\subset U^{\star}\setminus\{u\}$ does not satisfy Hall’s $d$-harem condition. If $X$ is not connected, then its neighbourhood is a disjoint union of the neighbourhoods of the connected components of $X$. Therefore without loss of generality we may assume that $X$ is connected. Since $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition, there exists $u^{\prime}\in X$ such that distance between $u^{\prime}$ and $u$ is equal to $2$. On the other hand the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $((U^{\star}\setminus\mathfrak{F}_{u})\cup\{u^{\perp}\})\cap\mathcal{B}(u)$. Therefore $X\not\subseteq((U^{\star}\setminus\{u\})\cup\{u^{\perp}\})\cap\mathcal{B}(u)$. It follows that there exists $u^{\prime\prime}\in X\setminus\mathcal{B}(u)$. The distance between $u^{\prime}$ and $u^{\prime\prime}$ is at least $14h(d(n+1))+$ 1, i.e. $|X|\geq h(d(n+1))+1$. By the definition of $\hat{h}$,
$|N_{\Gamma^{\star}}(X)|\geq d|X|+d.$
Thus
$|N_{\Gamma^{\star}\setminus\mathfrak{F}_{u}}(X)|\geq|N_{\Gamma^{\star}}(X)|-(d-1)\geq d|X|+1,$
i.e. $X$ satisfies Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$. ∎
By this claim there is $X\subset V^{\star}\setminus\mathfrak{F}_{u}$ that does not satisfy Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$. Take the smallest connected one. Observe that $u$ must be in $N_{\Gamma^{\star}}(X)$. Indeed, otherwise $|N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})}(X)|=|N_{\Gamma^{\star}}(X)|$. Since $\Gamma^{\star}$ satisfies $c.e.H.h.c.(d)$ we get a contradiction.
Claim 6.8.
For any vertex $v$ from $X$ the distance between $u$ and $v$ cannot exceed $12h(d(n+1))-$ 1.
Proof of Claim 6.8.
Since $\Gamma$ is bipartite, then distance between $u$ and any element of the set $X$ is odd. If for some $v\in X$ the distance between $u$ and $v$ is at least $12h(d(n+1))+$ 1, then $X$ has at least $h(d(n+1))$ elements. Since $\widehat{h}$ is a witness of $c.e.H.h.c.(d)$ for $\Gamma^{\star}$, we have
$|N_{\Gamma^{\star}\setminus\mathfrak{F}_{u}}(X)|\geq|N_{\Gamma^{\star}}(X)|-1\geq\frac{1}{d}|X|+1-1\geq\frac{1}{d}|X|,$
which contradicts that $X$ does not satisfy Hall’s $d$-harem condition. ∎
As a consequence of Claim 6.8, $X\subset(\mathcal{B}(u)\setminus\mathcal{S}(u))$.
Claim 6.9.
The element $u^{\perp}$ belongs to $N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}}(X)$ and the distance between $u^{\perp}$ and $u$ is at most $2h(d(n+1))$.
Proof of Claim 6.9.
Since $X\subset(\mathcal{B}(u)\setminus\mathcal{S}(u))$ and $\mathfrak{M}$ satisfies the conditions of the perfect $(1,d)$-matching for all subsets of $((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})\cap(\mathcal{B}(u)\setminus\mathcal{S}(u))$, then
$|N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}}(X)|\geq\frac{1}{d}|X|>|N_{\Gamma^{\star}\setminus\mathfrak{F}_{u}}(X)|.$
Clearly $u^{\perp}$ is in $N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}}(X)$.
If the distance between $u^{\perp}$ and $u$ is at least $22h(nd+1)+$ 2 then there exists an element of $X$ which is at distance at least $12h(nd+1)+$ 1 from $u$. Contradiction with Claim 6.8. ∎
We now see that $\{u,u^{\perp}\}\subset N_{\Gamma^{\star}\cup\mathfrak{S}}(X)$. Let us prove that Hall’s $d$-harem condition is satisfied in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for subsets of $(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$. By Claim 6.7 all subsets of $U^{\star}\setminus\{u\}$ satisfy Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$ and so in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ too. Therefore it remains to consider the case of subsets of $(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$ that contain $u^{\perp}$. Also note that the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ contained in $\mathcal{B}(u)\setminus\mathcal{S}(u)$. In particular trying to verify Hall’s $d$-harem condition in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for $Y\subset(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$ we may assume that $Y\nsubseteq\mathcal{B}(u)$. As above we also assume that $Y$ is connected.
Let $u_{1}\in Y\setminus\mathcal{B}(u)$. The distance between $u$ and $u_{1}$ at least $44h(d(n+1))+$ 4. Since $u^{\perp}\in N_{\Gamma^{\star}\cup\mathfrak{S}}(X)\cap Y$, we see by Claim 6.9 that there is some $u_{2}\in Y$ at the distance $\leq 2h(d(n+1))$ from $u$. Then the distance between $u_{1}$ and $u_{2}$ is at least $4h(d(n+1))+4-(2h(d(n+1))+2)$, i.e at least $22h(d(n+1))+$ 2. Thus $|Y|>h(d(n+1))$, and
$|N_{\Gamma}(Y)|-\frac{1}{d}|Y|\geq d(n+1).$
On the other hand,
$|N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}}(Y)|\geq|N_{\Gamma}(Y)|-n(d-1),$
hence:
$|N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}}(Y)|-\frac{1}{d}|Y|\geq|N_{\Gamma}(Y)|-\frac{1}{d}|Y|-n(d-1)\geq n\geq 0.$
It follows that Hall’s $d$-harem condition holds for any $Y\subset(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$ such that $u^{\perp}\in Y$. This finishes the case of subsets of $(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$.
It remains to show that Hall’s $d$-harem condition is satisfied in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for any $Y\subset V\cap((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})$. We may assume that $N_{\Gamma^{\star}\cup\mathfrak{S}}(Y)\cap\{u,u^{\perp}\}\neq\emptyset$. Indeed, otherwise apply the assumption that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition. Also note again, that the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ contained in $\mathcal{B}(u)\setminus\mathcal{S}(u)$. In particular we may assume that $Y\nsubseteq(\mathcal{B}(u)\setminus\mathcal{S}(u))$. As above we only consider connected $Y\subset(V\cap((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})$.
Take any element $v_{1}$ of $Y\setminus(\mathcal{B}(u)\setminus\mathcal{S}(u))$. It is clear that the distance from $u$ to $v_{1}$ is at least $34h(d(n+1))+$ 3. If $u\in N_{\Gamma^{\star}}(Y)$ then $Y$ contains a vertex at the distance $1$ from $u$. If $u\not\in N_{\Gamma^{\star}}(Y)$ then by Claim 6.9 $Y$ contains a vertex at the distance $12h(d(n+1))+$ 1 from $u$. In any case we can find $v_{2}\in Y$ at the distance at most $12h(d(n+1))+$ 1 from $u$. We see that the distance between $v_{1}$ and $v_{2}$ is at least $22h(d(n+1))+$ 2. Thus $Y$ has at least $h(d(n+1))$ elements and
$|N_{\Gamma(Y)}|-\frac{1}{d}|Y|\geq d(n+1).$
On the other hand using
$|N_{\Gamma^{\star}\cup\mathfrak{S}\setminus\mathfrak{F}_{u}}(Y)|\geq|N_{\Gamma(Y)}|-n$
we see that:
$|N_{\Gamma^{\star}\cup\mathfrak{S}\setminus\mathfrak{F}_{u}}(Y)|-\frac{1}{d}|Y|\geq|N_{\Gamma(Y)}|-n-\frac{1}{d}|Y|\geq d(n+1)-n=n(d-1)+d\geq 0,$
i.e. Hall’s $d$-harem condition is also satisfied. ∎
The last of technical lemmas concerns finding a witness of $c.e.H.h.c.(d)$ in a subgraph.
Lemma 6.10.
Assume that $\Gamma=(U,V,E)$ is a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$ with witness $h$. Let
  • $\mathfrak{M}$ be an $(1,d-1)$-matching in $\Gamma$ consisting of $\ell$ fans,
  • $\Gamma^{\star}=(U^{\star},V^{\star})$ be $\Gamma$ without $\mathfrak{M}$,
  • $\Gamma^{\star}$ satisfy Hall’s $d$-Harem condition.
Then
$\widehat{h}(m)=\left\{\begin{array}[]{rr}0,\quad\text{if}\quad m=0,\\ h(m+\ell d),\quad\text{if}\quad m>0.\end{array}\right.$
is a witness of $c.e.H.h.c.(d)$ for $\Gamma^{\star}$.
Proof.
Since $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition we may assume that $m>0$. First, we prove that $\widehat{h}$ is a witness of $c.e.H.h.c.(d)$ for subsets of $U^{\star}$. The inequality $|X|\geq\widehat{h}(m)$ implies $|X|\geq h(m+\ell d)$, i.e.
$|N_{\Gamma}(X)|-d|X|-\ell d\geq m.$
Since $d\geq 2$ and
$|N_{\Gamma^{\star}}(X)|\geq|N_{\Gamma}(X)|-\ell(d-1),$
we have
$|N_{\Gamma^{\star}}(X)|-d|X|\geq|N_{\Gamma}(X)|-\ell(d-1)-d|X|\geq m.$
In order to show that $\widehat{h}$ works for $V^{\star}$ we start with the observation that for each $X\subset V^{\star}$,
$|N_{\Gamma^{\star}}(X)|\geq|N_{\Gamma}(X)|-\ell.$
Therefore using $d\geq 2$ we see that when $|X|\geq\widehat{h}(m)$,
$|N_{\Gamma}(X)|-\frac{1}{d}|X|-m-\ell d\geq 0.$
It follows that:
$|N_{\Gamma^{\star}}(X)|-\frac{1}{d}|X|\geq|N_{\Gamma}(X)|-\frac{1}{d}|X|-\ell d\geq m.$
7. Graphs constructed in parts 1 and 2 of each step satisfy $c.e.H.h.c.(d)$
7.1. Claims 5.2, 5.6
Before stating the lemma proving claims 5.2, 5.6 we remind some of the relevant notation used at the $n+1$-st step of the construction.
  • $\Gamma^{(n)}$ is $U^{(n)}$-reflected;
  • $\Gamma^{(n)\star}$ is a subgraph of $\Gamma^{(n)}$ obtained by removal of $(1,d-1)$-fans with roots belonging to the set $U^{(n)\perp}$;
  • $\Gamma^{(n)\star}$ satisfies $c.e.H.h.c.(d)$;
  • $|U^{(n)\perp}|\leq n$;
  • $\mathfrak{M}^{1}_{n}$ is a finite $(1,d)$-matching in the bipartite graph $\mathcal{B}^{(n)\star}(u_{n})=\mathcal{B}^{(n)}(u_{n})\cap\Gamma^{(n)\star}$. It satisfies the conditions of the perfect $(1,d)$-matchings for all vertices that are at the distance less than $\max\{4h(3d(n+1))+3,5\}$ from $u_{n}$.
Lemma 7.1.
The bipartite graph $\Gamma^{(n)\star}(-u_{n})$ is highly computable for any $n$. Furthermore, one of the following holds:
  • $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$;
  • the graph $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$ for some vertex $u_{j}^{\perp}\in U^{(n)\perp}$.
In the latter case the element $u^{\perp}_{j}$ can be computed and the bipartite graph $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ is highly computable.
Remark 7.2.
If $U^{(n)\perp}=\emptyset$, then by Lemma 7.1 the graph $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$. In particular, Claim 5.2 holds.
Proof of Lemma 7.1.
Let
$\widehat{h}_{1}(m)=\left\{\begin{array}[]{rr}0,\quad\text{if}\quad m=0,\\ h(m+(3n+1)d),\quad\text{if}\quad m>0.\end{array}\right.$
and
$\widehat{h}_{2}(m)=\left\{\begin{array}[]{rr}0,\quad\text{if}\quad m=0,\\ h(m+3nd),\quad\text{if}\quad m>0.\end{array}\right.$
Claim 7.3.
Assume that one of the graphs $\Gamma^{(n)\star}(-u_{n})$ or $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions. Then in the first case the function $\widehat{h}_{1}$ witnesses $c.e.H.h.c.(d)$ for $\Gamma^{(n)\star}(-u_{n})$ and in the second case the function $\widehat{h}_{2}$ witnesses $c.e.H.h.c.(d)$ for $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$.
Proof.
Apply Lemma 6.10. In the first case view $\Gamma^{(0)}$ as $\Gamma$ and $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$ in this lemma. The number $\ell$ from the lemma is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by one, corresponding to the fan of $u_{n}$. Thus $\ell\leq 3n+1$. Since $\Gamma^{(n)\star}(-u_{n})$ satisfies Hall’s $d$-harem conditions, the lemma works.
In the second case view $\Gamma^{(0)}$ as $\Gamma$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ as $\Gamma^{\star}$. To compute the number corresponding to $\ell$ from the lemma we should take the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ together, then increase it by one, corresponding to the fan of $u_{n}$, and then subtract one, corresponding to the fan of $u_{j}^{\perp}$. Therefore $\ell\leq 3n$. Since $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. ∎
It is clear that both $\Gamma^{(n)\star}(-u_{n})$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ are highly computable bipartite graphs (assuming that $u^{\perp}_{j}$ is computed). Thus it is enough to show that one of these graphs satisfies Hall’s $d$-harem condition (then apply the claim). We know that the graph $\Gamma^{(n)\star}$ satisfies Hall’s $d$-harem condition. Let $\mathfrak{v}$ denote the only vertex from the set $\{v^{0}_{n,1},\ldots,v^{0}_{n,d}\}$ that belongs to $\Gamma^{(n)\star}(-u_{n})$. By Proposition 6.3 the choice of $u_{n},v^{0}_{n,1},\ldots,v^{0}_{n,d}$ ensures that $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$ satisfies Hall’s $d$-harem condition as well. Let $\mathfrak{M}$ denotes the perfect $(1,d)$-matching in $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$.
Since $U^{(n)\star}(-u_{n})=U^{(n)\star}(-u_{n},-\mathfrak{v})$, for $X\subset U^{(n)\star}(-u_{n})$ we have
(1) $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|\geq|N_{\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})}(X)|\geq d|X|.$
The corresponding property also holds for all subsets of $V^{(n)\star}(-u_{n})$ that do not contain $\mathfrak{v}$. Therefore if $\Gamma^{(n)\star}(-u_{n})$ does not satisfy Hall’s $d$-harem condition, then a witness of this is a finite $X\subset V^{(n)\star}(-u_{n})$ which contains $\mathfrak{v}$.
Claim 7.4.
If $X$ is a connected subset of $V^{(n)\star}(-u_{n})$ and $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$, then $X\subset(\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n}))$.
Proof.
Observe that by the choice of the radius of the ball $\mathcal{B}(u_{n})$, for any connected $X\nsubseteq(\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n}))$ with $\mathfrak{v}\in X$ we have $|X|\geq 2h(d(3n+1))+1$. Therefore
$|N_{\Gamma^{(0)}}(X)|-\frac{1}{d}|X|\geq d(3n+1).$
On the other hand at every step at most 3 fans from $\Gamma$ are added to the matching, hence $|U^{(0)}\setminus U^{(n)\star}(-u_{n})|\leq 3n+1$. As a consequence
$|N_{\Gamma^{(n)\star}(-u_{n})}(X)|\geq|N_{\Gamma^{(0)}}(X)|-(3n+1)\geq|N_{\Gamma^{(0)}}(X)|-d(3n+1)\geq\frac{1}{d}|X|.$
Since the neighbourhood of any set is decomposed into a disjoint union of the neighbourhoods of connected subsets, applying the claim we see that if $\Gamma^{(n)\star}(-u_{n})$ does not satisfy Hall’s $d$-harem condition, this is witnessed by a finite connected $X\subset V^{(n)\star}(-u_{n})\cap(\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n}))$ which contains $\mathfrak{v}$. Using high computability of $\Gamma^{(n)}$ and finiteness of $\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n})$, we compute $X$ with these properties.
We will now prove that there exists $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$. By the inequality $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$ it suffices to show that
(2) $|N_{\Gamma^{(n)}(-u_{n})}(X)|\geq\frac{1}{d}|X|.$
First, consider the case when $X=\{\mathfrak{v}\}$. The vertex $u_{\mathfrak{v}}$ either belongs to $U^{(n)\star}(-u_{n})$ or to $U^{(n)\perp}$. By inequality (1) and the fact that $U^{(n)\perp}$ consist of $(1,d-1)$-fans, in each case we have $|N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})|\geq d-1$. Moreover, the equality
(3) $|N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})|=d-1$
implies that $v_{u_{n}}\not\in\Gamma^{(n)}$. Indeed, if $v_{u_{n}}\in\Gamma^{(n)}$ then $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})$ by reflectedness. On the other hand equality (3) implies that $(u_{\mathfrak{v}},v_{u_{n}})\in\Gamma^{(n)\perp}$. Thus $(u_{n},\mathfrak{v})$ would be added to the matching $M$ at the first part of $n+1$-st step of the construction, i.e. $\mathfrak{v}\notin\Gamma^{(n)\star}(-u_{n})$, a contradiction. Indeed, if $v_{u_{n}}\in\Gamma^{(n)}$ then $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})$ (by reflectedness). On the other hand by the induction assumption, equality (3) implies that $(u_{\mathfrak{v}},v_{u_{n}})\in\Gamma^{(n)\perp}$. Thus $(u_{n},\mathfrak{v})$ would be added to the matching $M$ at the first part of $n+1$-st step of the construction, i.e. $\mathfrak{v}\notin\Gamma^{(n)\star}(-u_{n})$, a contradiction.
Our next observation is
$|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|\geq|N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})|-1.$
This follows by $U^{(n)}$-reflectedness of $\Gamma^{(n)}$: $v_{u_{n}}$ is the only possible element adjacent to $u_{\mathfrak{v}}$ that does not have the left copy in $U^{(n)}(-u_{n})$. Additionally note that the equality
(4) $|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|=|N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})|-1$
holds only if $v_{u_{n}}\in\Gamma^{(n)}(-u_{n})$.
Therefore equalities (3), (4) are not consistent, i.e.:
$|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|>(d-1)-1\mbox{ and }|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|\geq\frac{1}{d}.$
Since $X$ is a singleton,
$|N_{\Gamma^{(n)}(-u_{n})}(X)|\geq\frac{1}{d}|X|.$
As we mentioned above this means that there is some $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$. Using high computability of $\Gamma^{(n)}(-u_{n})$ we can compute it.
Then $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|\geq 1$, and we have $|X|>d$ by the assumption $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$. Consider this case. Let $U_{X}:=\{u\in U:v_{u}\in X\}$ in $\Gamma^{(n)}(-u_{n})$. Applying the fact that $V^{(n)}$ is a subset of the copy of $U^{(n)}$ we arrive at two possibilities:
  1. (i)
    $v_{u_{n}}\notin X$ and $|X|=|U_{X}|$;
  2. (ii)
    $v_{u_{n}}\in X$ and $|X|=|U_{X}|+1$.
In either case the inequality (2) follows from Lemma 6.2. Let $\mathfrak{S}$ denote the fan from the matching $\mathfrak{M}$ containing $v_{u_{n}}$ and $\mathfrak{S}^{\prime}$ denote $\mathfrak{S}$ with $v_{u_{n}}$ removed. The conditions of Lemma 6.2 are satisfied if we consider $\Gamma^{(n)}(-u_{n})$ as $\Gamma$ in that lemma, $v_{u_{n}}$ as $v$, and $\Gamma^{(n)\perp}(+\mathfrak{v})\cup\mathfrak{S}^{\prime}$ as $\Gamma^{\perp}$. Indeed, $\Gamma^{(n)}(-u_{n},-v_{u_{n}})$ is $U^{(n)}(-u_{n},-v_{u_{n}})$-reflected. Moreover, the corresponding graph $\Gamma^{\star}$ from the lemma is obtained by removal of $\Gamma^{(n)\perp}(+\mathfrak{v})\cup\mathfrak{S}^{\prime}$ from $\Gamma^{(n)}(-u_{n},-v_{u_{n}})$, therefore it is the same as $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$ with $\mathfrak{S}$ removed. Since $\mathfrak{S}$ is a fan from a perfect $(1,d)$- matching in $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$, we know that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition.
Therefore in case (i) by Lemma 6.2:
$|N_{\Gamma^{(n)}(-u_{n})}(X)|\geq(d-1)|X|-1.$
This fact combined with inequalities $d\geq 2$ and $|X|>d$ implies that
$|N_{\Gamma^{(n)}(-u_{n})}(X)|\geq\frac{1}{d}|X|.$
In case (ii) by Lemma 6.2:
$|N_{\Gamma^{(n)}(-u_{n})}(X\setminus\{v_{u_{n}}\})|\geq(d-1)(|X|-1)-1,$
[…]
Let us show that the inequality is strict. Indeed, by Lemma 6.2 the equality
$|N_{\Gamma^{(n)}(-u_{n})}(X\setminus\{v_{u_{n}}\})|=(d-1)(|X|-1)-1$
implies that $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(U_{X\setminus\{v_{u_{n}}\}})$ and $N_{\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})}(U_{X\setminus\{v_{u_{n}}\}})=\emptyset$. Therefore $v_{u_{n}}\in V^{(n)\perp}$. On the other hand, $v_{u_{n}}\in X\subset\Gamma^{(n)\star}$, a contradiction with the choice of $X$.
Again, inequalities $d\geq 2$ and $|X|>d$ implies
$|N_{\Gamma^{(n)}(-u_{n})}(X)|\geq(d-1)(|X|-1)\geq\frac{1}{d}|X|.$
Therefore, the assumption $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$ implies
$N_{\Gamma^{(n)}(-u_{n})}(X)\cap U^{(n)\perp}(-u_{n})\neq\emptyset,$
i.e. there exists some $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$. Then the vertex $u_{j}^{\perp}$ belongs to $N_{\Gamma^{(n)}(-u_{n})\cap\mathcal{B}(u_{n})}(X)$ (recall that $X\subseteq\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n})$) and we find it effectively. Let us denote by $v^{\perp}_{j,1},\ldots,v^{\perp}_{j,d-1}$ the $(d-1)$-tuple of vertices adjacent to $u_{j}^{\perp}$ in $\Gamma^{(n)\perp}$.
[…]
In order to prove that the graph $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem condition, we construct a perfect $(1,d)$-matching in the ball $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})\cap\mathcal{B}(u_{n})$. We set
$M^{\prime}:=(\mathfrak{M}^{1}_{n}\setminus\{(u_{n},v^{0}_{n,1}),\ldots,(u_{n},v^{0}_{n,d}),(u^{\prime}_{0},v^{\prime}_{0}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n-1})\})\cup$
$\{(u^{\perp}_{j},v^{\perp}_{j,1}),\ldots,(u^{\perp}_{j},v^{\perp}_{j,d-1}),(u^{\perp}_{j},v^{\prime}_{0}),(u^{\prime}_{0},v^{\prime}_{1}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n})\},$
where $v^{\prime}_{n}=\mathfrak{v}\in\{v^{0}_{n,1},\ldots,v^{0}_{n,d}\}$.
Figure 11. We replace the red fans in the matching $\mathfrak{M}^{1}_{n}$ by the blue fans to obtain the matching $M^{\prime}$ in $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$.
We have obtained $M^{\prime}$ by removing $d$ edges incident to $u_{n}$, adding $d$ edges incident to $u^{\perp}_{j}$, and the following replacement: for each of $u^{\prime}_{i}$ we replace one edge incident to it by another incident edge (then $\mathfrak{v}$ becomes adjacent to one edge in $M^{\prime}$). It follows that the matching $M^{\prime}$ satisfies the conditions of the perfect $(1,d)$-matchings for all susbsets of $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})\cap(\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n}))$.
To finish the proof notice that the conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}$ as $\Gamma^{\star}$, $M^{\prime}$ as $\mathfrak{M}$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u_{n}$ is considered as $u$ and $u^{\perp}_{j}$ as $u^{\perp}$. Therefore by Lemma 6.6, $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem condition. ∎
7.2. Notation used in proof of Lemma 7.5.
Before stating the second lemma, we remind the reader the relevant notation from the construction.
  • In Case 3A
    • $\mathfrak{U}:=U^{(n+1)}\setminus U^{(n)\perp}$
    • $\mathfrak{V}:=V^{(n+1)}\setminus V^{(n)\perp}$
    • $\mathfrak{T}:=(\mathfrak{U},\mathfrak{V},\mathfrak{E})$, where $\mathfrak{E}$ is induced in $\Gamma$ by the sets of vertices $\mathfrak{U},\mathfrak{V}$.
    • $\dot{U}^{(n)\perp}=U^{(n)\perp}\cap U^{(n+1)}$;
    • $\dot{V}^{(n)\perp}=V^{(n)\perp}\cap V^{(n+1)}$
    • $\dot{\Gamma}^{(n)\perp}:=(\dot{U}^{(n)\perp},\dot{V}^{(n)\perp},\dot{E}^{(n)\perp})$, where $\dot{E}^{(n)\perp}$ is defined according to the set of fans of elements from $\dot{U}^{(n)\perp}$ putted into $\dot{\Gamma}^{(n)\perp}$.
  • In Case 3B
    • $\mathfrak{U}:=U^{(n+1)}\setminus(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\})$
    • $\mathfrak{V}:=V^{(n+1)}\setminus(V^{(n)\perp}\cup\{\dot{v}_{n,i}^{\perp}:1\leq i\leq d-1\})$.
    • $\mathfrak{T}:=(\mathfrak{U},\mathfrak{V},\mathfrak{E})$, where $\mathfrak{E}$ is induced in $\Gamma$ by the sets of vertices $\mathfrak{U},\mathfrak{V}$.
    • $\dot{U}^{(n)\perp}:=(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\})\cap U^{(n+1)}$ with $\dot{V}^{(n)\perp}$ defined accordingly: $\dot{V}^{(n)\perp}:=(V^{(n)\perp}\cup\{\dot{v}_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}$;
    • $\dot{\Gamma}^{(n)\perp}:=(\dot{U}^{(n)\perp},\dot{V}^{(n)\perp},\dot{E}^{(n)\perp})$, where $\dot{E}^{(n)\perp}$ is defined according to the set of fans of elements from $\dot{U}^{(n)\perp}$ putted into $\dot{\Gamma}^{(n)\perp}$.
  • $\mathfrak{M}^{2}_{n}$ is a finite $(1,d)$-matching in the bipartite graph $\mathcal{B}^{(n)\star}(u^{l+1}_{n})=\mathcal{B}^{(n)}(u^{l+1}_{n})\cap\Gamma^{(n)\star}$. It satisfies the conditions of the perfect $(1,d)$-matchings for all vertices that are at the distance less than $\max\{4h(3d(n+1))+3,5\}$ from $u^{l+1}_{n}$;
  • $\dot{v}^{l+1}_{n,i}$, $1\leq i\leq d$ (or $d-1$), are vertices adjacent to $u^{l+1}_{n}$ in $\mathfrak{M}^{2}_{n}$ (or in $\Gamma^{(n)\perp}$);
  • $\dot{v}_{1},\dot{v}_{2}$ are the elements among $\dot{v}^{l+1}_{n,i}$, $1\leq i\leq d$ (or $d-1$), which are not added to $M^{l+1}_{n}$; there are at most two of them.
Lemma 7.5.
The bipartite graph $\mathfrak{T}$ is highly computable for any $n$. Moreover, one of the following holds:
  • $\mathfrak{T}$ satisfies $c.e.H.h.c.(d)$;
  • the graph $\mathfrak{T}(+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$ for some $u_{j}^{\perp}\in\dot{U}^{(n)\perp}$;
  • the graph $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$ for some $u_{i}^{\perp},u_{j}^{\perp}\in\dot{U}^{(n)\perp}$.
In the latter cases, the elements $u_{i}^{\perp},u_{j}^{\perp}$ can be computed and the corresponding bipartite graphs are highly computable.
Remark 7.6.
If $|\dot{U}^{(n)\perp}|\leq 1$, then Lemma 7.5 can be restated as follows. One of the following holds:
  • $\mathfrak{T}$ is a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$;
  • $\Gamma^{(n+1)}$ is a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$.
Therefore this Lemma proves Claim 5.4.
Proof of Lemma 7.5.
As in the proof of Lemma 7.1 we begin by showing that each of graphs from the formulation satisfies $c.e.H.h.c.(d)$ as long as it satisfies Hall’s $d$-harem condition.
Let
$\widehat{h}_{1}(m)=\left\{\begin{array}[]{rr}0,\quad\text{if}\quad m=0,\\ h(m+(3n+3)d),\quad\text{if}\quad m>0.\end{array}\right.,$
$\widehat{h}_{2}(m)=\left\{\begin{array}[]{rr}0,\quad\text{if}\quad m=0,\\ h(m+(3n+2)d),\quad\text{if}\quad m>0.\end{array}\right.$
and
$\widehat{h}_{3}(m)=\left\{\begin{array}[]{rr}0,\quad\text{if}\quad m=0,\\ h(m+(3n+1)d),\quad\text{if}\quad m>0.\end{array}\right.$
Claim 7.7.
Assume that one of the graphs $\mathfrak{T}$, or $\mathfrak{T}(+u_{j}^{\perp})$ or $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions. Then in the first case the function $\widehat{h}_{1}$ witnesses $c.e.H.h.c.(d)$ for $\mathfrak{T}$, in the second case the function $\widehat{h}_{2}$ witnesses $c.e.H.h.c.(d)$ for $\mathfrak{T}(+u_{j}^{\perp})$ and in the third case the function $\widehat{h}_{3}$ witnesses $c.e.H.h.c.(d)$ for $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$.
Proof.
It is a counterpart of Claim 7.3. In the first case view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}$ as $\Gamma^{\star}$ in Lemma 6.10, i.e. $\ell$ from this lemma is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by three, corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$. Thus $\ell\leq 3n+3$. Since $\mathfrak{T}$ satisfies Hall’s $d$-harem conditions, the lemma works.
In the second case view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}(+u_{j}^{\perp})$ as $\Gamma^{\star}$ i.e. $\ell$ from Lemma 6.10 is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by three, corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$ and decreased by one, corresponding to the fan of $u_{j}^{\perp}$. Therefore it is at most $23n+$ 2. Since $\mathfrak{T}(+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. In the third case view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ as $\Gamma^{\star}$ i.e. $\ell$ from this lemma is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by three, corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$ and decreased by two, corresponding to the fans of $u_{i}^{\perp},u_{j}^{\perp}$. Therefore it is at most $13n+$ 1. Since $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. ∎
Since (assuming that $u^{\perp}_{i},u^{\perp}_{j}$ are computed) each of the graphs from the claim is a highly computable bipartite graph, it is enough to show that at least one of them satisfies Hall’s $d$-harem condition (then apply the claim).
Assume that $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition. Let $u^{l+1}_{n}$ be the root of the last fan added to the matching $M_{n}$ in the second part of the $n$-th step; it belongs to the produced cycle of length $2$. Recall that $\dot{v}_{1},\dot{v}_{2}$ denote the vertices of the form $\dot{v}^{l+1}_{n,i}$ that were not added to $M^{l+1}_{n}$. In particular in Case 3A there is only one such a vertex and in Case 3B there is either one or two such vertices. From now on we consider the case of two additional vertices: $\dot{v}_{1}$ and $\dot{v}_{2}$. The case of only one of them, $\dot{v}_{1}$, is similar.
In the rest of the proof we will use the following two lemmas.
Lemma 7.8.
Depending on the existence of $\dot{v}_{2}$ either the graph $\mathfrak{T}(-\dot{v}_{1})$ or $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition.
Proof.
We only consider the case when $\dot{v}_{2}$ exists, the other one is analogous. It follows from the procedure that:
$\mathfrak{U}(-\dot{v}_{1},-\dot{v}_{2})=U^{(n)\star}(-u_{n})\setminus\{u^{l+1}_{n},\dot{u}^{\perp}_{n}\}$
and
$\mathfrak{V}(-\dot{v}_{1},-\dot{v}_{2})=V^{(n)\star}(-u_{n})\setminus\{\dot{v}^{l+1}_{n,1},\ldots,\dot{v}^{l+1}_{n,d},\dot{v}^{\perp}_{n,1},\ldots,\dot{v}^{\perp}_{n,d}\}.$
Since $\mathfrak{M}^{2}_{n}$ is a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u^{l+1}_{n})$ and $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ is obtained from $\Gamma^{(n)\star}(-u_{n})$ by removal of two fans from $\mathfrak{M}^{2}_{n}$, by Proposition 6.3 we know that $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. ∎
Lemma 7.9.
For any $X\subset V^{(n+1)}$, we have
$|N_{\Gamma^{(n+1)}}(X)|\geq\frac{1}{d}|X|.$
Proof.
The inequality follows from Lemma 6.1. Indeed, consider $\Gamma^{(n+1)}$ to be $\Gamma$ from that lemma. Note that the construction guarantees that $\Gamma^{(n+1)}$ is $U^{(n+1)}$-reflected. Depending on existence of $\dot{v}_{2}$ consider either $\dot{\Gamma}^{(n)\perp}(+\dot{v}_{1})$ or $\dot{\Gamma}^{(n)\perp}(+\dot{v}_{1},+\dot{v}_{2})$ to be $\Gamma^{\perp}$ from that lemma. Then the corresponding graph $\Gamma^{\star}$ from the lemma is equal to either $\mathfrak{T}(-\dot{v}_{1})$ or $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ and by Lemma 7.8 it satisfies Hall’s $d$-harem condition. Therefore conditions of Lemma 6.1 are satisfied. ∎
Applying Lemma 7.8 we see that the graph $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. Let $\mathfrak{M}$ denote the perfect $(1,d)$-matching in $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$.
It is clear that the inequality $|N_{\mathfrak{T}}(X)|\geq|N_{\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})}(X)|$ holds for all $X\subset\mathfrak{U}$. This inequality also holds for subsets of $\mathfrak{V}$ which do not intersect $\{\dot{v}_{1},\dot{v}_{2}\}$. Therefore if $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition then a witness of this is some finite subset of $\mathfrak{V}$ containing at least one of $\dot{v}_{1},\dot{v}_{2}$.
The rest of the proof is divided into two parts.
Part 1. We check whether the graph $\mathfrak{T}(-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. If it does, we set
  • $M^{\prime}$ is a matching obtained by Proposition 5.1 in $\mathfrak{T}(-\dot{v}_{2})\cap\mathcal{B}(u^{l+1}_{n})$;
  • $\dot{\Gamma}^{(n+1)\perp}:=\dot{\Gamma}^{(n)\perp}$;
  • $\Gamma^{\prime}=\mathfrak{T}$, denoting $\Gamma^{\prime}=(U^{\prime},V^{\prime})$,
and finish Part 1 of the proof.
If it does not, then there exists a minimal connected set $X$ such that $\dot{v}_{1}\in X\subset\mathfrak{V}(-\dot{v}_{2})$ and $|N_{\mathfrak{T}(-\dot{v}_{2})}(X)|<\frac{1}{d}|X|$. Analogously to the Claim 7.4, it can be showed that such $X$ is a subset of $(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$ and as a consequence we can compute $X$ with these properties.
Observe that $|N_{\Gamma^{(n+1)}}(X)|\geq\frac{1}{d}|X|$ by Lemma 7.9.
The inequalities
$|N_{\Gamma^{(n+1)}}(X)|\geq\frac{1}{d}|X|\,\mbox{ and }\,|N_{\mathfrak{T}(-\dot{v}_{2})}(X)|<\frac{1}{d}|X|$
imply
$N_{\Gamma^{(n+1)}}(X)\cap\dot{U}^{(n)\perp}\neq\emptyset,$
i.e. there exists some $u^{\perp}_{j}\in N_{\Gamma^{(n+1)}}(X)$. Since $\Gamma^{(n+1)}$ is highly computable, the element $u^{\perp}_{j}$ and its fan from $\dot{\Gamma}^{(n)\perp}$ can be computed. Similarly as in Lemma 7.1 we denote by $v^{\perp}_{j,1},\ldots,v^{\perp}_{j,d-1}$ the remaining vertices of the fan containing $u^{\perp}_{j}$.
[…]
We now apply an argument similar to one from the proof of Lemma 7.1. We set
$M^{\prime}:=\mathfrak{M}^{2}_{n}\setminus\{(u^{l+1}_{n},\dot{v}^{l+1}_{n,1}),\ldots,(u^{l+1}_{n},\dot{v}^{l+1}_{n,d}),(u^{\prime}_{0},v^{\prime}_{0}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n-1})\}\cup$
$\{(u^{\perp}_{j},v^{\perp}_{j,1}),\ldots,(u^{\perp}_{j},v^{\perp}_{j,d-1}),(u^{\perp}_{j},v^{\prime}_{0}),(u^{\prime}_{0},v^{\prime}_{1}),\ldots,(u^{\prime}_{n-1},v^{\prime}_{n})\},$
where $v^{\prime}_{n}=\dot{v}_{1}$. Observe that
$(\mathfrak{U}\cup\{u_{j}^{\perp}\},(\mathfrak{V}\cup\{v_{j,1}^{\perp},\ldots,v_{j,d-1}^{\perp}\})\setminus\{\dot{v}_{2}\})=\mathfrak{T}(+u_{j}^{\perp})\setminus\{\dot{v}_{2}\}=\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2}).$
We have obtained $M^{\prime}$ from $\mathfrak{M}^{2}_{n}$ by removing the $d$-fan of $u^{l+1}_{n}$, adding the $d$-fan of $u^{\perp}_{j}$, and the following replacement: for each of $u^{\prime}_{i}$ we replace one edge incident to it by another incident edge (then $\dot{v}_{1}$ becomes adjacent to one edge in $M^{\prime}$). Since $\mathfrak{M}^{2}_{n}$ was a $(1,d)$-perfect matching in the ball $\mathcal{B}(u^{l+1}_{n})\cap\Gamma^{(n)\star}$, it follows that matching $M^{\prime}$ satisfies the conditions of the perfect $(1,d)$-matchings for all susbsets of $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})\cap(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$. Therefore $M^{\prime}$ is a perfect $(1,d)$-matching in the ball $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})\cap\mathcal{B}(u^{l+1}_{n})$.
The conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$, $M^{\prime}$ as $\mathfrak{M}$ and graph $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u^{l+1}_{n}$ is considered as $u$ and $u^{\perp}_{j}$ is considered as $u^{\perp}$ in that lemma. Therefore by Lemma 6.6 graph $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition.
We define $\Gamma^{\prime}:=\mathfrak{T}(+u_{j}^{\perp})$, denote $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ and put
$\dot{\Gamma}^{(n+1)\perp}=(\dot{U}^{(n)\perp}\setminus\{u_{j}^{\perp}\},\dot{V}^{(n)\perp}\setminus\{v_{j,k}^{\perp}:1\leq k\leq d-1\}).$
This ends the first part of the proof.
Part 2. We check whether the graph $\Gamma^{\prime}$ satisfies Hall’s $d$-harem condition. If it does, then by Part 1 $\Gamma^{\prime}$ has to be equal to $\mathfrak{T}(+u_{j}^{\perp})$ and the proof is finished by the second option of the formulation.
If it does not then again, by repeating the reasoning of Part 1, we see that we can effectively recognize a minimal connected set $X$ such that $\dot{v}_{2}\in X\subset V^{\prime}\cap(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$ and $|N_{\Gamma^{\prime}}(X)|<\frac{1}{d}|X|$. Again, using Lemma 7.9 we obtain
$N_{\Gamma^{(n+1)}}(X)\cap\dot{U}^{(n)\perp}\neq\emptyset,$
i.e. there exists some $u^{\perp}_{i}\in N_{\Gamma^{(n+1)}}(X)$ and by high computability of $\Gamma^{(n+1)}$ it can be computed. As usual, we denote by $v^{\perp}_{i,1},\ldots,v^{\perp}_{i,d-1}$ the vertices adjacent to $u^{\perp}_{j}$ in $\dot{\Gamma}^{(n+1)\perp}$.
[…]
$M^{\prime\prime}:=M^{\prime}\setminus\{(u^{\prime}_{0},v^{\prime}_{0}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n-1})\}\cup$
$\{(u^{\perp}_{i},v^{\perp}_{i,1}),\ldots,(u^{\perp}_{i},v^{\perp}_{i,d-1}),(u^{\perp}_{i},v^{\prime}_{0}),(u^{\prime}_{0},v^{\prime}_{1}),\ldots,(u^{\prime}_{n-1},v^{\prime}_{n})\}.$
We have obtained $M^{\prime\prime}$ from $M^{\prime}$ by adding the $d$ edges incident to $u^{\perp}_{i}$, and replacing one edge in matching for each of $u^{\prime}_{k}$ in such a way, that $\dot{v}_{2}$is adjacent to one edge in $M^{\prime\prime}$. It follows that matching $M^{\prime\prime}$ is the perfect $(1,d)$-matching in the ball $\Gamma^{\prime}(+u_{i}^{\perp})\cap(\mathcal{B}(u^{l+1}_{n})$.
The conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$, $M^{\prime\prime}$ as $\mathfrak{M}$ and $\Gamma^{\prime}(+u_{i}^{\perp})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u^{l+1}_{n}$ is considered as $u$ and $u^{\perp}_{i}$ as $u^{\perp}$ in that lemma. Therefore by Lemma 6.6, $\Gamma^{\prime}(+u_{i}^{\perp})$ satisfies Hall’s $d$-harem condition.
The final argument depends on two possible outputs of Part 1. If the graph $\mathfrak{T}(-\dot{v}_{2})$ does not satisfy Hall’s $d$-harem condition (i.e. $u^{\perp}_{j}$ is involved), then $M^{\prime\prime}$ is a perfect $(1,d)$-matching in the graph $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$. If it does, we redefine $u^{\perp}_{j}:=u^{\perp}_{i}$ and then $M^{\prime\prime}$ becomes a perfect $(1,d)$-matching in the graph $\mathfrak{T}(+u_{j}^{\perp})$. Therefore if $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition, then either $\mathfrak{T}(+u_{j}^{\perp})$ or $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies this condition. ∎
8. Proof of the Main Theorem
Proof.
Let us apply the construction of Section 5. This construction works modulo Claims 5.2, 5.4, 5.6, 5.7. Claims 5.2 and 5.6 follow from Lemma 7.1. Claims 5.4 and 5.7 follow from Lemma 7.5. The union $\bigcup\limits_{i=1}^{n}M_{i}$ consist always only of $(1,d-1)$-fans, thus the final set of edges $M$ is a $(1,d-1)$-matching.
For every $u\in U$ there is a step where an edge incident to $u$ is added to $M$. Then if the copy $v_{u}$ was not added to $M$ earlier, in the second part of this step $v_{u}$ is also added to $M$. It follows that $M$ is a perfect $(1,d-1)$-matching of the graph $\Gamma$. This property also implies computability of this matching. Indeed, the construction guarantees that for any $n$ all edges from $M$ of the forms $(n,j)$ and $(k,n)$ belong to $\bigcup\limits_{i=1}^{n}M_{i}$. Therefore there is an algorithm which computes all such pairs from $M$ for any $n$. We see that the function $f$ realized by this matching is computable.
It remains to show that $f$ has controlled sizes of its cycle. Since the edges $(u_{0},v^{0}_{0,1})$ and $(v_{u_{0}},u_{v^{0}_{0,1}})$ are added to $M$ at step 1 we see $f^{2}(u_{0})=f(u_{v^{0}_{0,1}})=u_{0}$. Since $u_{0}=1$, condition $(i)$ of Definition 2.4 is satisfied.
Until the end of the proof we will abuse the notation and use a natural number $n$ both for vertex and number of step. Note that as a vertex $n$ appears in $M$ at the $n$-th step at latest. This follows from minimality of $u_{n}$ in $U^{(n-1)}$. The length of a cycle created at the $n$-th step cannot be greater than $\max\{2,n\}$ so if $n$ is in a cycle then $f^{i}(n)=n$ for some $i\leq n$ and condition $(ii)$ of Definition 2.4 is satisfied.
It remains to show that condition $(iii)$ is satisfied. Let $\mathfrak{f}_{i}$ be the partial function (living in $\mathbb{N}$) realized by $M_{i}$ and $f_{n}$ be the partial function realized by $\bigcup\limits_{i=0}^{n}M_{i}$. Let $F_{n}=\{(f_{n}(m),m):m\in\mathsf{Dom}(f_{n})\}$ be the graph of $f_{n}$ on $\mathsf{Dom}(f_{n})\cup\mathsf{Rng}(f_{n})$. Then $M_{n}=\{(\mathfrak{f}_{n}(m),m):m\in\mathsf{Dom}(\mathfrak{f}_{n})\}$ is the graph of $\mathfrak{f}_{n}$ and is a subgraph of $F_{n}$ too.
Observe that $F_{n}$ has at most $n+1$ connected components. Indeed, each subgraph $M_{i}$ has exactly one connected component. Furthermore, one of the following properties holds:
  • $M_{n}$ is a connected component of $F_{n}$;
  • $M_{n}$ is a fan with a root which already appears in $F_{n-1}$ as a vertex of degree $1$ (see the way how fans of $U$-elements are added to $M$).
The construction guarantees that $F_{n}$ consists only of vertices of degree $1$ and $d$. When a vertex has degree $1$, then its $F_{n}$-neighbourhood is exactly its $f_{n}$-image. When a vertex has degree $d$, then its $f_{n}$-image and $d-1$ preimages are in $F_{n}$. In particular each connected component of $F_{n}$ contains a cycle. The length of the cycle is not greater than $\max\{n,2\}$.
Since the value $f_{n-1}(n)$ is defined, $n$ belongs to some connected component of $f_{n-1}$. Thus there exist $k$ and $l$ such that $f_{n-1}^{k+l}(n)=f_{n-1}^{k}(n)$. These $k$ and $l$ work for the equality $f^{k+l}(n)=f^{k}(n)$. It remains to show that $k$ can be bounded by $2n$ and $l$ by $n$.
The latter estimate is easy: the biggest cycle that can be constructed before the $n$-th step does not have more than $\max\{n-1,2\}$ elements. Below we will bound $l$ by $n$ for simplicity.
In order to show that $k$ is bounded by $2n$ let us estimate the size of a subset of $U$ without a cycle that can be added to the matching $M$ in the process of the $n$-th step of the construction. It must consist of elements of $U^{(n-1)\perp}$ added to the matching at the iteration of part $2$ of the $n$-th step together with $u_{n}$. Therefore we can bound it by the maximal possible number of elements in $U^{(n-1)\perp}$ increased by $1$.
Let us denote the number of elements from $U^{(s)\perp}$ added to $M$ at the $s$-th step by $\ell_{s}$. Clearly, $f^{\ell_{s}+1}(n)\in M_{j}$ for some $s\leq n$ and $j\leq n-1$. If no cycle is constructed in the $j$-th step then $f^{\ell_{j}+1}(f^{\ell_{s}+1}(n))\in M_{i}$ for some $i\leq j-1$. Iterating this argument we arrive at
$k\leq\sum\limits_{s=1}^{n}(\ell_{s}+1).$
Since at the $n$-th step the size of $\bigcup\limits_{s=1}^{n}U^{(s)\perp}$ does not exceed $n-1$,
$\sum\limits_{s=1}^{n}\ell_{s}\leq n-1.$
We see that $k\leq 2n-1$. Thus condition $(iii)$ of Definition 2.4 is satisfied. ∎
9. Schneider’s theorem and computable entourages of coarse spaces
We are ready to prove Theorem A from the introduction. We repeat it.
Theorem 9.1.
Let $d\geq 3$. Let a coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that for any finite $F\subseteq\mathbb{N}$ we have $|E[F]|\geq(d+2)|F|$. Then there exists a computable $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a $d$-regular forest. Moreover, there exist an algorithm which for any $m,n\in\mathbb{N}$ recognizes if $m$ and $n$ are in the same connected component of $\Gamma(E^{\prime})$.
Proof.
Let $E$ be a highly computable symmetric entourage as in the formulation. Consider the graph $\Gamma(R)$ defined for the symmetric relation $R:=E\setminus\Delta_{\mathbb{N}}\subseteq\mathbb{N}\times\mathbb{N}$. Since the coarse space $(\mathbb{N},\mathcal{E})$ is a bounded geometry, the neighbourhood of each vertex is finite. The graph $\Gamma(R)$ is highly computable by high computability of $E$.
We are going to apply Theorem 4.1 to $\Gamma(R)$. It remains to show that $\Gamma(R)$ satisfies $c.e.H.h.c.(d)$. To see this note:
$|R[F]|\geq|E[F]\setminus F|\geq|E[F]|-|F|\geq(d+1)|F|.$
Thus, for all finite sets $X\subset\mathbb{N}$:
$|N_{\Gamma(R)}(X)|-d|X|\geq(d+1)|X|-d|X|=|X|.$
In particular the inequality $n\leq|X|$ implies $n\leq|N(X)|-d|X|\leq|N(X)|-\frac{1}{d}|X|$. Since the identity map on $\mathbb{N}$ is a total computable function, it follows that $\Gamma(R)$ satisfies $c.e.H.h.c.(d)$.
Let us apply Theorem 4.1. We obtain a computable function $f$ realizing a perfect $(1,d-1)$ matching in the graph $\Gamma(R)$ with controlled sizes of its cycles. In order to construct $E^{\prime}\in\mathcal{E}$ as in the formulation we adapt the proof of Theorem 2.2 of [Sch18].
Since $f$ is a total, computable, surjective, $(d-1)$ to $1$ function, the graph of $f$ (denoted by $\Gamma(f)$) is computable and $d$-regular. We remind the reader that $f$ satisfies the following properties:
  1. (i)
    $f^{2}(1)=1$;
  2. (ii)
    if $n\geq 2$ and $f^{i}(n)=n$ then $i\leq n$;
  3. (iii)
    if $n\geq 2$ and for all $i\leq n$ we have $f^{i}(n)\neq n$ then there exist $k\leq 2n$ and $l\leq n$ such that $f^{k+l}(n)=f^{k}(n)$;
  4. (iv)
    for each $n$ the pair $(n,f(n))$ belongs to $R$.
Since $R=E\setminus\Delta_{X}$ the last property implies that $f$ does not have fixed points. Let
$P(f)=\{n\in\mathbb{N}\,|\,\exists m\geq 1(f^{m}(n)=n)\}\,\mbox{ and }\,P_{0}(f)=\{n\in P(f)\,|\,\forall m\geq 1(f^{m}(n)\geq n)\},$
i.e. the union of all cycles and the set of minimal representatives of cycles. By property (ii) there is an algorithm which for every $n\in\mathbb{N}$ verifies whether $n\in P(f)$, i.e. $P(f)$ is computable. Observe that $P_{0}(f)$ is computable too. Indeed, $1$ obviously belongs to $P_{0}(f)$. When $n\geq 2$ and $n\in P(f)$, then verifying if $f^{i}(n)\geq n$ for all $i\leq n$ we can check whether $n\in P_{0}(f)$ (apply (ii) again).
Since there is no component with two disjoint cycles we see that whenever $n,m\in P_{0}(f)$ and $n\neq m$, then $n$ and $m$ do not belong to the same connected component of the graph $\Gamma(f)$. Thus
$P(f)=\dot{\bigsqcup}_{n\in P_{0}(f)}\{f^{m}(n)\,|\,m\in\mathbb{N}\}.$
There is an algorithm which for every $n\in\mathbb{N}$ finds the $P_{0}(f)$-representative of the connected component of $n$. Indeed, if for example $n\notin P(f)$ then applying (iii) we can find $i\leq 2n$ such that $f^{i}(n)\in P(f)$ and later $j\leq n$ such that $f^{i+j}(n)\in P_{0}(f)$.
Based on this we want to construct a new computable function $f_{\star}$ such that its graph (denoted by $\Gamma(f_{\star})$) is a computable $d$-regular forest. Let us start with two auxiliary functions $g,h:P_{0}(f)\times\mathbb{N}\rightarrow\mathbb{N}$ such that, for all $n\in P_{0}(f)$ and $m\geq 1$ the following properties hold:
  • $g(n,0)=n$ and $h(n,0)=f(n)$;
  • $\{g(n,m),h(n,m)\}\cap P(f)=\emptyset$;
  • $f(g(n,m))=g(n,m-1)$, and $f(h(n,m))=h(n,m-1)$.
We want $g,h$ to be computable functions. Since $P_{0}(f)$ and $P(f)$ are computable and the graph $\Gamma(f)$ is computable and $d$-regular the following rule gives required algorithm. Given $n\in P_{0}(f)$ and $m\geq 1$ and having defined $g(n,m-1)$ find the minimal $x$ such that $x\notin P(f)$ and $f(x)=g(n,m-1)$. Then let $g(n,m)=x$. The definition of $h(n,m)$ is similar. Clearly $g$ and $h$ are injective and have disjoint ranges.
Now we define $f_{\star}:\mathbb{N}\rightarrow\mathbb{N}$ for $x\in\mathbb{N}$ in the following way:
$f_{\star}(x)=\left\{\begin{array}[]{ll}g(y,m+2)&\text{ if }x=g(y,m)\text{ for }y\in P_{0}(f)\text{ and even }m\geq 0,\\ g(y,m-2)&\text{ if }x=g(y,m)\text{ for }y\in P_{0}(f)\text{ and odd }m\geq 3,\\ f^{2}(x)&\text{ if }x=h(y,m)\text{ for }y\in P_{0}(f)\text{ and }m\geq 2,\\ f(x)&\text{ otherwise.}\end{array}\right.$
We remind the reader that for each $n$ there is $i\leq 3n$ such that $f^{i}(n)\in P_{0}(f)$. Thus for any $m^{\prime}>3n$ the number $n$ does not belong to $\{g(y,m^{\prime}),h(y,m^{\prime})\}$. This guarantees that $f_{\star}$ is computable. Facts that $\Gamma(f_{\star})\in\mathcal{E}$, the function $f_{\star}$ does not have cycles and for each $x\in\mathbb{N}$ the size $|f^{-1}_{\star}(x)|$ is $d-1$ are proved in [Sch18, proof of Theorem 2.2]. Therefore the graph $\Gamma(f_{\star})$ is a computable $d$-regular forest.
To see the last statement of the theorem note that if $C\subset\mathbb{N}$ is a connected components of $\Gamma(f)$ then it is also the set of vertices of a connected component of $\Gamma(f_{\star})$. Since for each $n$ one can compute $n^{\prime}\in P_{0}(f)$ such that $n$ and $n^{\prime}$ are in same connected component of $\Gamma(f)$ we have an algorithm which for any $n$ and $m$ verifies whether $n$ and $m$ are in the same tree in $\Gamma(f_{\star})$. ∎
We now translate Theorem A into a statement about wobbling groups. The following theorem is Theorem B from the introduction. It is a computable version of Corollary 2.3 of [Sch18].
Theorem 9.2.
Let the coarse space $(\mathbb{N},\mathcal{E})$ of a bounded geometry be non-amenable. Furthermore, assume that there exists a highly computable symmetric $E\in\mathcal{E}$ such that $|E[F]|\geq 6|F|$ for all finite $F\subset\mathbb{N}$. Then there are two computable permutations $\sigma,\pi\in Sym(\mathbb{N})$ such that $\langle\sigma,\pi\rangle$ is a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$.
Proof.
By Theorem A there is $E^{\prime}\in\mathcal{E}$ such that $\Gamma(E^{\prime})$ is a computable $4$-regular forest. Moreover, there is an algorithm which for each $n,m\in\mathbb{N}$ recognizes if $n$ and $m$ are in the same connected component of $\Gamma(E^{\prime})$. Further, there is another algorithm which or each $n$ finds four natural numbers $n_{1}<n_{2}<n_{3}<n_{4}$ such that either $(n,n_{i})\in E^{\prime}$ or $(n_{i},n)\in E^{\prime}$. Since $\Gamma(E^{\prime})$ is a computable $4$-regular forest, for each $n,m\in\mathbb{N}$, we can compute $B_{m}(n)\subset\mathbb{N}$, the $m$-ball of $n$, by a uniform algorithm.
The standard Cayley graph of the free group on two generators is isomorphic to the $4$-regular tree so it is clear that there exist permutations $\sigma$,$\pi$ as in the formulation. We will show that there exist computable ones. The construction is by induction. At every step we construct two finite partial isometries of $\Gamma(E^{\prime})$.
We will use the following notion. We say that a partial permutation $\sigma$ is loxodromic on a finite, connected set $D$, if:
  • it preserves distance i.e. for all $n,m\in D$ we have $d(n,m)=d(\sigma(n),\sigma(m))$;
  • there is a maximal line in the set $D\cup\sigma(D)$ i.e. there exists $d_{0},\ldots,d_{k}\in D\cup\sigma(D)$ for $k=\max\{d(a,b)|a,b\in D\}+1$, such that $\sigma(d_{0})=d_{1},\ldots,\sigma(d_{k-1})=d_{k}$ where $d_{0}\in D\setminus\sigma(D)$ and $d_{k}\in\sigma(D)\setminus D$.
At the first step of the construction of $\sigma$ and $\pi$ we find $1_{1},1_{2},1_{3},1_{4}$ and set $\sigma_{1}(1)=1_{1},\sigma_{1}(1_{2})=1,\pi_{1}(1)=1_{3},\pi_{1}(1_{4})=1$. Let $D_{1}=B_{1}(1)=\{1,1_{1},1_{2},1_{3},1_{4}\}$. We extend $\sigma_{1}$ and $\pi_{1}$ to loxodromic maps on $D_{1}$ such that $\sigma_{1}(D_{1})=B_{1}(\sigma_{1}(1))$ and $\pi_{1}(D_{1})=B_{1}(\pi_{1}(1))$.
Assume that before the $n$-th step of the construction two finite partial permutations $\sigma_{n-1}$ and $\pi_{n-1})$ are defined so that they are loxodromic maps on some set $D_{n-1}$. For all $a\in D_{n-1}$ we set $\sigma_{n}(a)=\sigma_{n-1}(a)$ and $\pi_{n}(a)=\pi_{n-1}(a)$. Let $a_{n}$ be the first number in $\mathbb{N}\setminus D_{n-1}$. Let $X$ be the connected component of $\Gamma(E^{\prime})$ containing $a_{n}$. If it does not meet $D_{n-1}$ then we find the corresponding $(a_{n})_{1},(a_{n})_{2},(a_{n})_{3},(a_{n})_{4}$, and define $\sigma_{n}$ and $\pi_{n}$ on $\{a_{n},(a_{n})_{1},(a_{n})_{2},(a_{n})_{3},(a_{n})_{4}\}$ as at the first step. Then put $D_{n}=D_{n-1}\cup B_{1}(a_{n})$.
It is worth mentioning a this stage that according the assumptions on $\Gamma(E^{\prime})$ the question if $X\cap D_{n-1}\not=\emptyset$ is decidable: just check if $a_{n}$ is in the same component with elements of $D_{n-1}$. Now assume that the answer is positive. Then find the minimal $i<n$ such that $a_{i}\in X$ and the minimal $m$ such that $a_{n}\in B_{m}(a_{i})$. Then we extend $\sigma_{n}$ and $\pi_{n}$ to a loxodromic partial permutations on $D_{n}=D_{n-1}\cup B_{m}(a_{i})$ in such way that for any $a\in D_{n}$ and any $i,j\in\{-1,1\}$ we have $\sigma^{i}_{n}(a)\neq\pi^{j}_{n}(a)$.
Let $\sigma=\lim\limits_{n\in\mathbb{N}}\sigma_{n}$, and $\pi=\lim\limits_{n\in\mathbb{N}}\pi_{n}$. These permutations act on each connected component of $\Gamma(E^{\prime})$ by translations along some line, with lines corresponding to $\sigma$ and $\pi$ having one common vertex which is the first $n$ that belongs to component. Clearly there are no points which are fixed by non-identity of $\langle\sigma,\pi\rangle$ and this group is are a free semi-regular subgroup of $\mathcal{W}(\mathbb{N},\mathcal{E})$. Permutations $\sigma$ and $\pi$ are computable by the construction.